### v1

# Multilevel polynomial partitioning and semialgebraic hypergraphs: regularity, Turán, and Zarankiewicz results

Abstract.
We prove three results about semialgebraic hypergraphs. First, we prove an optimal and oblivious regularity lemma. Fox–Pach–Suk proved that the class of $k$-uniform semialgebraic hypergraphs satisfies a very strong regularity lemma where the vertex set can be partitioned into $\mathrm{poly}(1/\varepsilon)$ pieces so that all but an $\varepsilon$-fraction of $k$-tuples of parts are homogeneous (either complete or empty). Our result improves the number of parts in the partition to $O_{d,k}((D/\varepsilon)^{d})$ where $d$ is the dimension of the ambient space and $D$ is a measure of the complexity of the hypergraph; additionally, the partition is oblivious to the edge set of the hypergraph. We give examples that show that the dependence on both $\varepsilon$ and $D$ is optimal.
From this regularity lemma we deduce the best-known Turán-type result for semialgebraic hypergraphs. We show that any such hypergraph with $N$ vertices and $\varepsilon N^{k}$ edges contains a complete $k$-uniform $k$-partite subhypergraph with $\Omega_{d,k}(D^{-d(k-1)}\varepsilon^{d(k-1)+1}N^{k})$ edges.
Third, we prove a Zarankiewicz-type result for semialgebraic hypergraphs. Previously, Fox–Pach–Sheffer–Suk–Zahl showed that a $K_{u,u}$-free semialgebraic graph on $N$ vertices has at most $O_{d,D,u}(N^{2d/(d+1)+o(1)})$ edges and Do extended this result to $K_{u,\ldots,u}^{(k)}$-free semialgebraic hypergraphs. We improve upon both of these results by removing the $o(1)$ in the exponent and making the dependence on $D$ and $u$ explicit and polynomial.
All three of these results follow from a novel “multilevel polynomial partitioning scheme” that efficiently partitions a point set $P\subset\mathbb{R}^{d}$ by low-complexity semialgebraic pieces. We prove this result using the polynomial method over varieties as developed by Walsh which extends the real polynomial partitioning technique of Guth–Katz.
Tidor was supported by a Stanford Science Fellowship
1. Introduction
Semialgebraic graphs and hypergraphs appear ubiquitously in discrete geometry and beyond. These are graphs whose vertex set is a set of points in $\mathbb{R}^{d}$ and whose edges are defined by a finite number of polynomial inequalities. They can represent incidences between points and arbitrary varieties, crossings between curves, and many other configurations arising in discrete geometry.
Definition 1.1.
A semialgebraic hypergraph in $\mathbb{R}^{d}$ is a $k$-uniform $k$-partite hypergraph $H=(P_{1}\sqcup\cdots\sqcup P_{k},E)$ where $P_{1},\ldots,P_{k}$ are each finite subsets of $\mathbb{R}^{d}$ such that the following holds. There are polynomials $g_{1},\ldots,g_{t}\in\mathbb{R}[x_{11},\ldots,x_{kd}]$ and a function $\Phi\colon\{-1,0,1\}^{t}\to\{0,1\}$ so that $((x_{11},\ldots,x_{1d}),\ldots,(x_{k1},\ldots,x_{kd}))\in P_{1}\times\cdots\times P_{k}$ is an edge if and only if
$\Phi(\operatorname{sgn}(g_{1}(x_{11},\ldots,x_{kd})),\ldots,\operatorname{sgn}(g_{t}(x_{11},\ldots,x_{kd})))=1.$
We call $\sum_{r=1}^{t}\deg g_{r}$ the total degree of $H$.
Remark.
Previous works use slightly different definitions of the complexity of a hypergraph. For example, [5] gives a semialgebraic hypergraph a complexity $(t,D)$ where $t$ is the number of polynomials used and $D$ is an upper bound on the degree of each polynomial as a function of $x_{i1},\ldots,x_{id}$ for each $i\in[k]$. A semialgebraic hypergraph of complexity $(t,D)$ thus has total degree at most $tkD$.
In this paper, we prove three main results about semialgebraic hypergraphs. The first concerns the regularity lemma for semialgebraic hypergraphs.
Definition 1.2.
For a $k$-partite $k$-uniform hypergraph $H=(P_{1}\sqcup\cdots\sqcup P_{k},E)$ we say that partitions $\Pi_{1},\ldots,\Pi_{k}$ of $P_{1},\ldots,P_{k}$ respectively form a homogeneous partition with error $\varepsilon$ if
$\sum\frac{|\pi_{1}|\cdots|\pi_{k}|}{|P_{1}|\cdots|P_{k}|}\leq\varepsilon$
where the sum is over $k$-tuples of parts $(\pi_{1},\ldots,\pi_{k})\in\Pi_{1}\times\cdots\times\Pi_{k}$ that are not homogeneous, i.e., such that $E[\pi_{1}\times\cdots\times\pi_{k}]$ is neither complete nor empty.
Fox, Pach, and Suk [5] proved that every semialgebraic hypergraph has a homogeneous partition with $O(\varepsilon^{-C})$ parts where $C$ only depends on the uniformity, dimension of the ambient space, and complexity of the hypergraph. Our main theorem gives the optimal dependence of the number of parts on the error $\varepsilon$ and the total degree $D$. Furthermore, the partition produced will be essentially oblivious to the hypergraph itself, only depending on the vertex sets as well as $\varepsilon$ and $D$.
Theorem 1.3.
Fix $d,k,D\geq 1$ and $\varepsilon>0$. Let $P_{1},\ldots,P_{k}\subset\mathbb{R}^{d}$ be finite sets. There exist partitions $\Pi_{1},\ldots,\Pi_{k}$ of $P_{1},\ldots,P_{k}$ respectively of size
$|\Pi_{1}|,\ldots,|\Pi_{k}|=O_{d,k}\left(\left(D/\varepsilon\right)^{d}\right).$
These have the property that for any semialgebraic hypergraph $H$ on vertex sets $P_{1},\ldots,P_{k}$ with total degree at most $D$, the partitions $\Pi_{1},\ldots,\Pi_{k}$ give a homogeneous partition of $H$ with error $\varepsilon$.
In Section 6, we give a family of examples showing that this bound on the number of parts is optimal. In other words, there exist semialgebraic hypergraphs for which any homogeneous partition with error $\varepsilon$ must have $\Omega_{d,k}((D/\varepsilon)^{d})$ parts. While this paper was in preparation, Rubin [8] proved a version of this result where the number of parts is bounded by $O_{d,k,D}((1/\varepsilon)^{d+o(1)})$. For a comparison of his techniques and ours, see Section 1.1.
A partition is called equitable if every pair of parts differs in size by at most 1. Via standard techniques one can deduce an equitable version of the main theorem with a small loss in the bound. In dimension $d=1$, our techniques give an equitable partition with no loss in the bound.
Corollary 1.4.
Fix $d,k,D\geq 1$ and $\varepsilon>0$. Let $P_{1},\ldots,P_{k}\subset\mathbb{R}^{d}$ be finite sets. There exist equitable partitions $\Pi_{1},\ldots,\Pi_{k}$ of $P_{1},\ldots,P_{k}$ respectively of size
$|\Pi_{1}|,\ldots,|\Pi_{k}|=O_{d,k}\left(D^{d}\varepsilon^{-(d+1)}\right).$
These have the property that for any semialgebraic hypergraph $H$ on vertex sets $P_{1},\ldots,P_{k}$ with total degree at most $D$, the partitions $\Pi_{1},\ldots,\Pi_{k}$ give a homogeneous partition of $H$ with error $\varepsilon$. Furthermore, if $d=1$, we can instead guarantee that
$|\Pi_{1}|,\ldots,|\Pi_{k}|=O_{k}\left(D/\varepsilon\right).$
One straightforward consequence of Theorem 1.3 is a Turán-type result. This improves upon a recent result of Rubin [9, Theorem 1.6] who proves the same $\varepsilon$-dependence without an explicit $D$-dependence.
Corollary 1.5.
Let $H=(P_{1}\sqcup\cdots\sqcup P_{k},E)$ be a semialgebraic hypergraph in $\mathbb{R}^{d}$ with total degree $D$ and at least $\varepsilon|P_{1}|\cdots|P_{k}|$ edges. Then there exist sets $S_{1}\subseteq P_{1},\ldots,S_{k}\subseteq P_{k}$ such that $S_{1}\times\cdots\times S_{k}\subseteq E$ and $|S_{1}|\cdots|S_{k}|=\Omega_{d,k}(D^{-d(k-1)}\varepsilon^{d(k-1)+1}|P_{1}|\cdots|P_{k}|)$. In fact, for $11\leq i\leq k-$ 1,
$|S_{i}|=\Omega_{d,k}\left((\varepsilon/D)^{d}|P_{i}|\right)$
and
$|S_{k}|=\Omega\left(\varepsilon|P_{k}|\right).$
Furthermore, our techniques also imply a Zarankiewicz result for semialgebraic hypergraphs. This asks for the maximum number of edges in an $N$-vertex $K_{u,\ldots,u}^{(k)}$-free semialgebraic hypergraph. The best previous result for semialgebraic graphs is due to Fox, Pach, Sheffer, Suk, and Zahl [4] and for semialgebraic hypergraphs is due to Do [3]. We improve upon their results in two ways; first, we remove the $N^{o(1)}$-term in their $N$-dependence and second, we make the dependence on $u$ and the total degree of the hypergraph explicit and polynomial.
Theorem 1.6.
Let $H=(P_{1}\sqcup\cdots\sqcup P_{k},E)$ be a semialgebraic hypergraph in $\mathbb{R}^{d}$ with total degree $D$. Suppose $|P_{1}|=\cdots=|P_{k}|=N$. If $H$ is $K^{(k)}_{u,\ldots,u}$-free, then the number of edges of $H$ is at most
$|E|=O_{d,k}\left(\sum_{i=1}^{k}u^{\frac{i}{(i-1)d+1}}D^{\frac{(i-1)(i+2)d}{2((i-1)d+1)}}N^{k-\frac{i}{(i-1)d+1}}\right).$
Later in the paper we will state and prove this result when $P_{1},\ldots,P_{k}$ have different sizes. We also remark that this and all previous results can easily be generalized to the case when $P_{1},\ldots,P_{k}$ live in $\mathbb{R}^{d_{1}},\ldots,\mathbb{R}^{d_{k}}$ for different dimensions $d_{1},\ldots,d_{k}$.
1.1. Proof ideas and comparison to previous work
Our results are proved using what we call a “multilevel polynomial partitioning scheme”. We construct a partition of a point set $P\subset\mathbb{R}^{d}$ as follows. Using the real polynomial partitioning technique first developed by Guth and Katz [6], we find a low degree polynomial $f$ that cuts $\mathbb{R}^{d}\setminus Z(f)$ into a number of open cells each with few points from $P$. For each open cell $C$ (i.e., connected component of $\mathbb{R}^{d}\setminus Z(f)$) we add $P\cap C$ as a part of our partition. It then remains to partition $P\cap Z(f)$. To handle these points we use polynomial partitioning again inside the variety $Z(f)$. We iterate this a total of $d$ times, reducing the dimension of the variety by one each iteration. To produce an efficient quantitative dependence, it is important that we use the method developed by Walsh [10] for real polynomial partitioning inside a given variety. Our multilevel polynomial partitioning scheme is qualitatively similar to previous work of Matoušek and Patáková [7] though our work has significantly stronger quantitative dependence.
To prove the semialgebraic regularity lemma (Theorem 1.3) we apply this partitioning technique to the vertex sets $P_{1},\ldots,P_{k}$ with a parameter $A$ chosen in terms of $d,k,\varepsilon,D$. It then remains to prove that this gives a homogeneous partition of any semialgebraic hypergraph $H$ on vertex set $P_{1}\sqcup\cdots\sqcup P_{k}$ with total degree at most $D$. To do this we use an argument that is inspired by recent work of Bukh and Vasileuski [2] on the same-type lemma. We view our $k$ partitions of $\mathbb{R}^{d}$ as giving a single partition of $\mathbb{R}^{dk}$. Then a $k$-tuple of cells is non-homogeneous only if it is properly crossed by one of the defining polynomials $g_{1},\ldots,g_{t}$ of the hypergraph $H$. We bound the number of cells crossed by the zero set of a polynomial using a Milnor–Thom-type result proved by Barone and Basu [1].
While this paper was in preparation, Rubin [8] independently proved a version of Theorem 1.3 where the number of parts is bounded by $O_{d,k,D}((1/\varepsilon)^{d+o(1)})$. His regularity lemma is also oblivious to the edge set of the hypergraph and has the additional advantage that the partition can be computed by an algorithm that runs in near linear time. However, his result has the disadvantage that the dependence of the number of parts on $D$ is not stated explicitly and is likely quite large, and the bound loses a factor of $\varepsilon^{-o(1)}$. Rubin also constructs his partitions via iterated applications of polynomial partitioning though the technical details differ quite a lot from this paper. In addition, he proves the correctness of this partition using an analysis in $\mathbb{R}^{dk}$ that also is inspired by the recent work of Bukh and Vasileuski [2].
Our Turán result (Corollary 1.5) is a straightforward corollary of the semialgebraic regularity lemma. To prove our Zarankiewicz result (Theorem 1.6) we combine the Turán result with a second application of the multilevel polynomial partitioning scheme.
Notation
We use standard graph theory notation. We interchangeably write $A=O(B)$ and $A\lesssim B$ to mean that there exists some absolute constant $C$ so that $A\leq CB$.
Acknowledgements
We thank Jacob Fox for numerous enlightening discussions as well as Huy Tuan Pham for helpful conversations about Section 6.
2. Multilevel polynomial partitioning
In this paper we use the word variety to refer to a (not-necessarily irreducible) algebraic variety in $\mathbb{C}^{d}$. For a variety $V$ we use the typical notion of degree and dimension. We write $V(\mathbb{R})$ for the real points of $V$, i.e., $V(\mathbb{R})=V\cap\mathbb{R}^{d}$. Given polynomials $Q_{1},\ldots,Q_{r}$, we write $Z(Q_{1},\ldots,Q_{r})$ for the variety they cut out.
Our main result is the following multilevel polynomial partitioning scheme. The input is a finite set of points $P\subset\mathbb{R}^{d}$ and a parameter $A\geq 1$. We produce a partition of $P$ into $O_{d}(A^{d})$ parts where each part of the partition is contained in a semialgebraic set $C$. Here $C$ is a connected component of $Z(Q_{1},\ldots,Q_{r})(\mathbb{R})\setminus Z(f)$ where $Q_{1},\ldots,Q_{r},f$ are polynomials of degree $O_{d}(A)$.
Theorem 2.1 (Multilevel polynomial partitioning).
Let $P\subset\mathbb{R}^{d}$ be a finite set of points. Given a parameter $A\geq 1$ we produce sets
$\mathcal{V}=\mathcal{V}_{0}\sqcup\mathcal{V}_{1}\sqcup\cdots\sqcup\mathcal{V}_{d}.$
For $0\leq i\leq d$, each element $V\in\mathcal{V}_{i}$ is a tuple
$V=(Q^{(V)},f_{V},P_{V})\in\mathcal{V}_{i}$
where
  • $Q^{(V)}=(Q^{(V)}_{1},\ldots,Q^{(V)}_{d-i})$ is a $(d-i)$-tuple of complex polynomials each of degree $O_{d}(A)$;
  • $f_{V}$ is a real polynomial of degree $O_{d}(A)$ that vanishes identically on every irreducible component of $Z(Q^{(V)})$ of dimension at least $i+1$; and
  • $P_{V}\subseteq Z(Q^{(V)})\cap P$ is a finite set of points.
These satisfy
$\sum_{V\in\mathcal{V}_{i}}\prod_{j=1}^{d-i}\deg Q^{(V)}_{j}\lesssim_{d}A^{d-i}$ (1)
for each $0\leq i\leq d$ as well as
$\sum_{V\in\mathcal{V}}|P_{V}|\lesssim_{d}|P|.$ (2)
For each $V\in\mathcal{V}$, define $\mathcal{C}_{V}$ to be the set of connected components of $Z(Q^{(V)})(\mathbb{R})\setminus Z(f_{V})$. Now define the set of cells as
$\mathcal{C}=\bigsqcup_{V\in\mathcal{V}}\mathcal{C}_{V}.$
For each $C\in\mathcal{C}_{V}$, define $P_{C}=C\cap P_{V}$. Then we have a partition
$P=\bigsqcup_{C\in\mathcal{C}}P_{C}$ (3)
satisfying
$|P_{C}|\lesssim_{d}\frac{|P_{V}|}{A^{i}\prod_{j=1}^{d-i}\deg Q_{j}^{(V)}}$ (4)
for each $0\leq i\leq d$, each $V\in\mathcal{V}_{i}$, and each $C\in\mathcal{C}_{V}$.
We defer the proof until the next section. For now we demonstrate how to use this multilevel polynomial partitioning scheme by using it to prove Theorem 1.3. This requires bounding the number of cells in the partition and bounding the number of $k$-tuples of bad cells. We deduce this from a Milnor–Thom-type result proved by Barone and Basu [1]. The deduction from Barone and Basu’s work is somewhat technical and proceeds in a similar way to how Walsh deduces [10, Theorem 1.5].
We place Barone and Basu’s result in the following form which we find more convenient to work with. For a variety $V\subseteq\mathbb{C}^{d}$ and a point $x\in V$, we write $\dim_{x}^{\mathbb{C}}(V)$ for the maximum dimension of an irreducible component of $V$ that contains $x$.
Theorem 2.2 ([1], see also [10]).
Given polynomials $Q_{1},\ldots,Q_{d-r}\in\mathbb{C}[x_{1},\ldots,x_{d}]$, the number of connected components of $Z(Q_{1},\ldots,Q_{d-r})(\mathbb{R})$ that contain a point $x$ with $\dim_{x}^{\mathbb{C}}(Z(Q_{1},\ldots,Q_{d-r}))=r$ is at most
$\lesssim_{d}\deg Q_{1}\cdots\deg Q_{d-r}\left(\max_{j=1,\ldots,{d-r}}\deg Q_{j}\right)^{r}.$
Note that by Krull’s theorem, every $x\in Z(Q_{1},\ldots,Q_{d-r})$ satisfies $\dim_{x}^{\mathbb{C}}(Z(Q_{1},\ldots,Q_{d-r}))\geq r$ so this result counts connected components whose dimension is as small as possible. The above result immediately from [10, Proposition 7.1] and [10, Lemma 7.2]. For completeness, we include this brief deduction in the appendix.
Now we are in a position to bound the number of cells produced by the multilevel polynomial partitioning scheme.
Lemma 2.3.
Given $A\geq 1$ and $P\subset\mathbb{R}^{d}$, let $V\in\mathcal{V}_{i}$ and $\mathcal{C}_{V}$ be as in Theorem 2.1 for some $0\leq i\leq d$. Then
$|\mathcal{C}_{V}|\lesssim_{d}A^{i}\prod_{j=1}^{d-i}\deg Q_{j}^{(V)}.$
Together with 1, this implies that $|\mathcal{C}|\lesssim_{d}A^{d}$.
Proof.
Consider a connected component $C$ of $Z(Q^{(V)})(\mathbb{R})\backslash Z(f_{V})$. Suppose $C$ is contained in $C^{\prime}$, a connected component of $Z(Q^{(V)})(\mathbb{R})$. If $C\neq C^{\prime}$, then $Z(f_{V})$ cuts $C^{\prime}$ and so $C$ “touches” $Z(f_{V})$. Thus we are in one of the three following cases:
  1. (i)
    $C$ is a connected component of $Z(Q^{(V)})(\mathbb{R})$; or
  2. (ii)
    there exists $\varepsilon_{C}>0$ such that $f_{V}$ takes all values $(0,\varepsilon_{C}]$ on $C$; or
  3. (iii)
    there exists $\varepsilon_{C}>0$ such that $f_{V}$ takes all values $[-\varepsilon_{C},0)$ on $C$.
Recall that $f_{V}$ vanishes on every irreducible component of $Z(Q^{(V)})$ of dimension at least $i+1$. Thus every $x\in C$ satisfies $\dim_{x}^{\mathbb{C}}(Z(Q^{(V)}))=i$. Therefore, by Theorem 2.2, the number of connected components $C$ in case (i) above is at most
$\lesssim_{d}\prod_{j=1}^{d-i}\deg Q_{j}^{(V)}\left(\max_{j}\deg Q_{j}^{(V)}\right)^{i}\lesssim_{d}A^{i}\prod_{j=1}^{d-i}\deg Q_{j}^{(V)}.$
To deal with the connected components in cases (ii) and (iii), pick $\varepsilon>0$ so that $\varepsilon<\varepsilon_{C}$ for all such $C$ and such that neither $f_{V}-\varepsilon$ nor $f_{V}+\varepsilon$ vanishes identically on any irreducible component of $Z(Q^{(V)})$. We can do this since there are only a finite number of $\varepsilon_{C}$ and also a finite number of bad $\varepsilon$ for which $f_{V}\pm\varepsilon$ vanishes identically on some irreducible component of $Z(Q^{(V)})$.
Then each $C$ in case (ii) contains a connected component of $Z(Q^{(V)})(\mathbb{R})\cap Z(f_{V}-\varepsilon)$. Moreover, for any $x\in Z(Q^{(V)},f_{V}-\varepsilon)$, we know that every irreducible component of $Z(Q^{(V)})$ which contains $x$ has dimension $i$. By our choice of $\varepsilon$ we know that every irreducible component of $Z(Q^{(V)},f_{V}-\varepsilon)$ containing $x$ has dimension $i-1$. In other words $\dim_{x}^{\mathbb{C}}(Z(Q^{(V)},f_{V}-\varepsilon))=i-1$. Therefore, applying Theorem 2.2 to $Z(Q^{(V)},f_{V}-\varepsilon)(\mathbb{R})$, we conclude that the number of connected components $C$ in case (ii) above is at most
$\lesssim_{d}\deg f_{V}\prod_{j=1}^{d-i}\deg Q_{j}^{(V)}\left(\max\left\{\deg f_{V},\deg Q_{1}^{(V)},\ldots,\deg Q_{d-i}^{(V)}\right\}\right)^{i-1}\lesssim_{d}A^{i}\prod_{j=1}^{d-i}\deg Q_{j}^{(V)}.$
Applying the same argument to $Z(Q^{(V)},f_{V}+\varepsilon)(\mathbb{R})$ gives the same bound for case (iii), completing the proof. ∎
By a similar but more involved argument we can prove the following lemma which is crucial for bounding the number of non-homogeneous $k$-tuples of cells.
Lemma 2.4.
Given finite sets $P_{1},\ldots,P_{k}\subset\mathbb{R}^{d}$, let $\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(k)}$ be the data produced by applying by Theorem 2.1 with parameter $A\geq 1$. Consider $V_{1}\in\mathcal{V}^{(1)}_{i_{1}},\ldots,V_{k}\in\mathcal{V}^{(k)}_{i_{k}}$ and a real polynomial $g\in\mathbb{R}[x_{11},\ldots,x_{kd}]$. The number of $k$-tuples of cells $(C_{1},\ldots,C_{k})\in\mathcal{C}_{V_{1}}\times\cdots\times\mathcal{C}_{V_{k}}$ such that $\operatorname{sgn}(g(x_{11},\ldots,x_{kd}))$ is not constant on $C_{1}\times\cdots\times C_{k}\subseteq\mathbb{R}^{dk}$ is at most
$\lesssim_{d,k}\deg g\cdot A^{i_{1}+\cdots+i_{k}-1}\prod_{\ell=1}^{k}\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{\ell})}.$
Proof.
If $\deg g\geq A$, then using Lemma 2.3 to bound $|\mathcal{C}_{V_{1}}\times\cdots\times\mathcal{C}_{V_{k}}|$ suffices to prove the desired result. Thus we assume that $\deg g\leq A$ from now on.
Define the list of polynomials
$\tilde{Q}=\bigsqcup_{\ell=1}^{k}\left\{Q_{1}^{(V_{\ell})}(x_{\ell 1},\ldots,x_{\ell d}),\ldots,Q_{d-i_{\ell}}^{(V_{\ell})}(x_{\ell 1},\ldots,x_{\ell d})\right\}.$
Next, define the real polynomial
$\tilde{f}=\prod_{\ell=1}^{k}f_{V_{\ell}}(x_{\ell 1},\ldots,x_{\ell d}).$
Note that $Z(\tilde{Q})(\mathbb{R})\setminus Z(\tilde{f})=\prod_{j=1}^{k}Z(Q^{(V_{j})})(\mathbb{R})\setminus Z(f_{V_{j}})$, so the set of connected components of $Z(\tilde{Q})(\mathbb{R})\setminus Z(\tilde{f})$ is exactly $\left\{C_{1}\times\cdots\times C_{k}:(C_{1},\ldots,C_{k})\in\mathcal{C}_{V_{1}}\times\cdots\times\mathcal{C}_{V_{k}}\right\}$.
Since $f_{V_{j}}$ vanishes identically on the all irreducible components of $Z(Q^{(V_{j})})$ with dimension at least $i_{j}+1$, we see that for each point $x\in Z(\tilde{Q})(\mathbb{R})\setminus Z(\tilde{f})$ satisfies $\dim_{x}^{\mathbb{C}}(Z(\tilde{Q}))=i_{1}+\cdots+i_{k}$.
Define $\mathcal{C}_{+}$ to be the set of connected components of $Z(\tilde{Q})(\mathbb{R})\setminus Z(\tilde{f})$ on which $g$ takes both positive values and 0. Similarly define $\mathcal{C}_{-}$ as the set of connected components on which $g$ takes both negative values and 0. Since $g\colon\mathbb{R}^{dk}\to\mathbb{R}$ is continuous in the Euclidean topology, it is clear that if $\operatorname{sgn}(g(x_{11},\ldots,x_{kd}))$ is not constant on a connected component $C$, then $C\in\mathcal{C}_{+}\cup\mathcal{C}_{-}$. We now bound $|\mathcal{C}_{+}|$; bounding $|\mathcal{C}_{-}|$ is symmetric.
Since $Z(\tilde{Q})$ has finitely many irreducible components, there are a finite number of bad $\delta\in\mathbb{R}$ such that $g-\delta$ vanishes identically on an irreducible component of $Z(\tilde{Q})$. For each $C\in\mathcal{C}_{+}$, there exists $\delta_{C}>0$ so that $g$ takes all values $[0,\delta_{C}]$ on $C$. Now pick $\delta>0$ so that $\delta<\delta_{C}$ for all $C\in\mathcal{C}_{+}$ and so that $g-\delta$ does not vanish identically on any irreducible component of $Z(\tilde{Q})$. Clearly each $C\in\mathcal{C}_{+}$ contains a connected component of $Z(\tilde{Q},g-\delta)(\mathbb{R})\setminus Z(\tilde{f})$.
Furthermore, since each point $x\in Z(\tilde{Q})(\mathbb{R})\setminus Z(\tilde{f})$ satisfies $\dim_{x}^{\mathbb{C}}(Z(\tilde{Q}))=i_{1}+\cdots+i_{k}$, by our choice of $\delta$, we see that $\dim_{x}^{\mathbb{C}}(Z(\tilde{Q},g-\delta))=i_{1}+\cdots+i_{k}-1$ for all $x\in Z(\tilde{Q},g-\delta)(\mathbb{R})\setminus Z(\tilde{f})$.
As in the proof of Lemma 2.3, each connected component $C$ of $Z(\tilde{Q},g-\delta)(\mathbb{R})\setminus Z(\tilde{f})$ satisfies one of the three following properties:
  1. (i)
    $C$ is a connected component of $Z(\tilde{Q},g-\delta)(\mathbb{R})$; or
  2. (ii)
    there exists $\varepsilon_{C}>0$ such that $\tilde{f}$ takes all values $(0,\varepsilon_{C}]$ on $C$; or
  3. (iii)
    there exists $\varepsilon_{C}>0$ such that $\tilde{f}$ takes all values $[-\varepsilon_{C},0)$ on $C$.
Pick $\varepsilon>0$ such that $\varepsilon<\varepsilon_{C}$ for all such $C$ and such that neither $\tilde{f}-\varepsilon$ nor $\tilde{f}+\varepsilon$ vanishes identically on any irreducible component of $Z(\tilde{Q},g-\delta)$. Then we see that for each point $x\in Z(\tilde{Q},g-\delta,\tilde{f}\pm\varepsilon)(\mathbb{R})\setminus Z(\tilde{f})$ satisfies $\dim_{x}^{\mathbb{C}}(Z(\tilde{Q},g-\delta,\tilde{f}\pm\varepsilon))=i_{1}+\cdots+i_{k}-2$.
Thus by Theorem 2.2, the number of $C$ in case (i) is bounded by
$\begin{split}\lesssim_{d,k}\deg g\prod_{\ell=1}^{k}\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{\ell})}&\left(\max\left\{\deg g,\deg Q_{1}^{(V_{1})},\ldots,\deg Q_{d-i_{k}}^{(V_{k})}\right\}\right)^{i_{1}+\cdots+i_{k}-1}\\ &\lesssim_{d,k}\deg g\cdot A^{i_{1}+\cdots+i_{k}-1}\prod_{\ell=1}^{k}\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{\ell})}.\end{split}$
Furthermore, each $C$ in case (ii) above contains a connected component of $Z(\tilde{Q},g-\delta,\tilde{f}-\varepsilon)(\mathbb{R})$ and each $C$ in case (iii) above contains a connected component of $Z(\tilde{Q},g-\delta,\tilde{f}+\varepsilon)(\mathbb{R})$. Again by Theorem 2.2, the number of $C$ in cases (ii) and (iii) is bounded by
$\begin{split}\lesssim_{d,k}\deg g\cdot\deg\tilde{f}\prod_{\ell=1}^{k}\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{\ell})}&\left(\max\left\{\deg g,\deg\tilde{f},\deg Q_{1}^{(V_{1})},\ldots,\deg Q_{d-i_{k}}^{(V_{k})}\right\}\right)^{i_{1}+\cdots+i_{k}-2}\\ &\lesssim_{d,k}\deg g\cdot A^{i_{1}+\cdots+i_{k}-1}\prod_{\ell=1}^{k}\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{\ell})}.\end{split}$
Summing these three cases gives the desired bound on $|\mathcal{C}_{+}|$. An identical argument bounds $|\mathcal{C}_{-}|$. ∎
Proof of Theorem 1.3.
For some large constant $C_{d,k}$, define $A=C_{d,k}D/\varepsilon\geq 1$. Let $\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(k)}$ be the data produced by applying Theorem 2.1 to $P_{1},\ldots,P_{k}$ respectively with parameter $A$. Let $\Pi_{1},\ldots,\Pi_{k}$ be the partitions of $P_{1},\ldots,P_{k}$ produced. Then by Lemma 2.3 and 1,
$|\Pi_{\ell}|=|\mathcal{C}^{(\ell)}|=\sum_{i=0}^{d}\sum_{V\in\mathcal{V}_{i}^{(\ell)}}|\mathcal{C}_{V}|\lesssim_{d}\sum_{i=0}^{d}\sum_{V\in\mathcal{V}_{i}^{(\ell)}}A^{i}\prod_{j=1}^{d-i}\deg Q_{j}^{(V)}\lesssim_{d}\sum_{i=0}^{d}A^{d}\lesssim A^{d}$
for each $1\leq\ell\leq k$.
Now let $H$ be any semialgebraic hypergraph on $P_{1},\ldots,P_{k}$ defined by polynomials $g_{1},\ldots,g_{t}$ with $\sum_{r}\deg g_{r}\leq D$. By Lemma 2.4, for each $g_{r}$ and each $V_{1}\in\mathcal{V}_{i_{1}}^{(1)},\ldots,V_{k}\in\mathcal{V}_{i_{k}}^{(k)}$, the number of $k$-tuples of cells in $\mathcal{C}_{V_{1}}\times\cdots\times\mathcal{C}_{V_{k}}$ on which $\operatorname{sgn}(g_{r}(x_{11},\ldots,x_{kd}))$ is not constant is bounded by
$\lesssim_{d,k}\deg g_{r}\cdot A^{i_{1}+\cdots+i_{k}-1}\prod_{\ell=1}^{k}\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{\ell})}.$
By 4, the fraction of $k$-tuples of vertices contributed by each $k$-tuple of cells $(C_{1},\ldots,C_{k})\in\mathcal{C}_{V_{1}}\times\cdots\times\mathcal{C}_{V_{k}}$ is
$\prod_{\ell=1}^{k}\frac{|P_{C_{\ell}}|}{|P_{\ell}|}\lesssim_{d,k}\prod_{\ell=1}^{k}\frac{|P_{V_{\ell}}|}{A^{i_{\ell}}|P_{\ell}|\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{\ell})}}.$
Summing over the contribution of a $k$-tuple of cells with this particular $g_{r},V_{1},\ldots,V_{k}$, we get a contribution of
$\lesssim_{d,k}\frac{\deg g_{r}}{A}\prod_{\ell=1}^{k}\frac{|P_{V_{\ell}}|}{|P_{\ell}|}.$
By 2, summing over all choices of $V_{1},\ldots,V_{k}$ gives a contribution of
$\lesssim_{d,k}\frac{\deg g_{r}}{A}.$
Finally, summing over the choices of $g_{r}$, we see that at most a
$\lesssim_{d,k}\sum_{r=1}^{t}\frac{\deg g_{r}}{A}\leq\frac{D}{A}=\frac{\varepsilon}{C_{d,k}}$
fraction of $k$-tuples of vertices lie in a $k$-tuple of cells on which $\operatorname{sgn}(g_{r}(x_{11},\ldots,x_{kd}))$ is not constant for some $g_{r}$. Choosing $C_{d,k}$ large enough makes the fraction at most $\varepsilon$, which means that $\Pi_{1},\ldots,\Pi_{k}$ form a homogeneous partition of $H$ with error $\varepsilon$. ∎
Proof of Corollary 1.4.
The deduction of the $d\geq 2$ case of Corollary 1.4 from Theorem 1.3 is standard (see, e.g., [5, Proof of Theorem 1.3]). We include a proof for completeness.
Let $\Pi_{1},\ldots,\Pi_{k}$ be the partitions of $P_{1},\ldots,P_{k}$ produced by Theorem 1.3 with parameters $d,k,D\geq 1$ and $\varepsilon/2$. For each $1\leq i\leq k$, say say $N_{i}=|P_{i}|$ and define $K_{i}=\left\lceil 4k\varepsilon^{-1}|\Pi_{i}|\right\rceil$. Index $P_{i}=\{p_{1},\ldots,p_{N_{i}}\}$ so that each part $\pi\in\Pi_{i}$ is an interval in this indexing. Now define $\Pi_{i}^{\prime}$ to be an equitable partition of $P_{i}$ into $K_{i}$ pieces, each of which is an interval of length $\left\lfloor N_{i}/K_{i}\right\rfloor$ or $\left\lceil N_{i}/K_{i}\right\rceil$. We claim that $\Pi_{1}^{\prime},\ldots,\Pi_{k}^{\prime}$ have the desired properties.
Call a part $\pi\in\Pi_{i}^{\prime}$ bad if it is not contained in a part of $\Pi_{i}$. Note that the number of bad parts is at most $|\Pi_{i}|$ and each bad part has size at most $2N_{i}/K_{i}$. For a semialgebraic hypergraph $H$ on $P_{1}\sqcup\cdots\sqcup P_{k}$, a $k$-tuple of vertices $(x_{1},\ldots,x_{k})$ is contained in a $k$-tuple of parts of $(\pi_{1},\ldots,\pi_{k})\in\Pi_{1}^{\prime}\times\cdots\times\Pi_{k}^{\prime}$ that is not homogeneous only if either it is contained in a $k$-tuple of parts of $\Pi_{1}\times\cdots\times\Pi_{k}$ that is not homogeneous or one of the $\pi_{i}$ is bad. The number of $k$-tuples of vertices of the first type is at most $(\varepsilon/2)N_{1}\cdots N_{k}$. The number of $k$-tuples of vertices of the second type is at most
$\sum_{i=1}^{k}\left(\prod_{j\neq i}N_{j}\right)\frac{2N_{i}}{K_{i}}|\Pi_{i}|=N_{1}\cdots N_{k}\sum_{i=1}^{k}\frac{2|\Pi_{i}|}{K_{i}}\leq\frac{\varepsilon}{2}N_{1}\cdots N_{k}.$
Therefore $\Pi_{1}^{\prime},\ldots,\Pi_{k}^{\prime}$ form a homogeneous partition of $H$ with error $\varepsilon$. Note that
$|\Pi_{i}^{\prime}|=K_{i}=\left\lceil 4k\varepsilon^{-1}|\Pi_{i}|\right\rceil=O_{d,k}(D^{d}\varepsilon^{-(d+1)}).$
Now we show how to save a factor of $\varepsilon$ when $d=1$. For $P_{i}\subset\mathbb{R}$ and any positive integer $A$, we can find $y_{1},\ldots,y_{A-1}$ so that
$P_{i}\cap(-\infty,y_{1}),P_{i}\cap(y_{1},y_{2}),\ldots,P_{i}\cap(y_{A-2},y_{A-1}),P_{i}\cap(y_{A-1},+\infty)$
is an equitable partition of $P_{i}$ into $A$ pieces. Instead of applying Theorem 2.1, we will construct the data that the theorem produces by hand. Define $\mathcal{V}_{0}=\emptyset$ and $\mathcal{V}_{1}=\left\{Q^{(\mathbb{C})},f_{\mathbb{C}},P_{\mathbb{C}}\right\}$ where $Q^{(\mathbb{C})}=\emptyset$, where $P_{\mathbb{C}}=P_{i}$, and where $f_{\mathbb{C}}(x)=\prod_{j=1}^{A-1}(x-y_{j})$. One can easily check that these satisfy all of the desired properties and that the partition of $P_{i}$ produced is the above equitable partition.
Applying this procedure for $i=1,\ldots,k$ with $A=C_{1,k}D/\varepsilon$, we produce equitable partitions $\Pi_{1},\ldots,\Pi_{k}$ of $P_{1},\ldots,P_{k}$. Then the proof of Theorem 1.3 shows that these partitions have the desired property. ∎
3. Constructing the partition
In this section, we explain how to construct the multilevel polynomial partitioning scheme in Theorem 2.1. The key tool is a quantitatively efficient real polynomial partitioning result over irreducible varieties introduced by Walsh [10]. This allows us to afford to do real polynomial partitioning with polynomials of super-constant degree and lets us avoid the $o(1)$-loss in the exponent. Other discrete geometry results have also been established by similar techniques in a later paper of Walsh [11]. The argument there is considerably more involved due to numerous technical difficulties in that setting.
To introduce Walsh’s real polynomial partitioning result, we need to define a more refined notion of degree that gives more precise quantitative control over a variety. This will be used extensively throughout our proof.
Definition 3.1 (Partial degrees).
For an irreducible variety $V\subseteq\mathbb{C}^{d}$ and $1\leq i\leq d-\dim V$, define the $i$-th partial degree, $\delta_{i}(V)$, to be the minimum positive integer such that there exist polynomials $f_{1},\ldots,f_{t}\in\mathbb{C}[x_{1},\ldots,x_{d}]$ of degree at most $\delta_{i}(V)$ such that $V\subseteq Z(f_{1},\ldots,f_{t})$ and such that every irreducible component of $Z(f_{1},\ldots,f_{t})$ containing $V$ has dimension at least $d-i$. We define $\delta_{0}(V)=0$ and $\delta_{i}(V)=\infty$ for $i>d-\dim V$.
By a short argument ([10, Lemma 2.3]) these are known to satisfy $\delta_{1}(V)\leq\delta_{2}(V)\leq\cdots\leq\delta_{d-\dim V}(V)$.
To give some intuition for this definition, consider polynomials $Q_{1},\ldots,Q_{d-r}\in\mathbb{C}[x_{1},\ldots x_{d}]$ satisfying $\deg Q_{1}\leq\cdots\leq\deg Q_{d-r}$. Suppose that $V=Z(Q_{1},\ldots,Q_{d-r})$ is irreducible. If $Z(Q_{1}),\ldots,Z(Q_{d-r})$ intersect nicely enough, (i.e., if $V$ is a complete intersection) then $V$ will have dimension $r$ and degree $\deg Q_{1}\cdots\deg Q_{d-r}$. Furthermore, in this case $\delta_{i}(V)=\deg Q_{i}$ for all $i$. Although not every irreducible variety is a complete intersection, Walsh’s techniques say that we are always approximately in this situation. In particular, for any irreducible variety $V\subseteq\mathbb{C}^{d}$ of dimension $r$, there exist polynomials $Q_{1},\ldots,Q_{d-r}\in\mathbb{C}[x_{1},\ldots,x_{d}]$ so that $V$ is an irreducible component of $Z(Q_{1},\ldots,Q_{d-r})$ and so that $\delta_{i}(V)=\deg Q_{i}$. Furthermore, $\deg V$ and $\deg Q_{1}\cdots\deg Q_{d-r}$ will be equal up to a multiplicative constant that only depends on the ambient dimension. We state the second part of this result below and something a little stronger than the first result as Proposition 3.5.
Proposition 3.2 ([10, Corollaries 5.4 and 5.7]).
For an irreducible variety $V\subseteq\mathbb{C}^{d}$,
$\prod_{i=1}^{d-\dim V}\delta_{i}(V)\lesssim_{d}\deg V\leq\prod_{i=1}^{d-\dim V}\delta_{i}(V).$
Having introduced the definition of partial degrees, we can now state the result of Walsh [10] that allows us to do efficient real polynomial partitioning over any irreducible varieties. The formulation below is slightly different from what is stated in his paper. We include a short deduction in the appendix.
Theorem 3.3 ([10, Theorem 3.2]).
Let $V\subseteq\mathbb{C}^{d}$ be an irreducible variety and let $P\subseteq V(\mathbb{R})$ be a finite set of points. For $A\geq 1$, there exists a polynomial $f\in\mathbb{R}[x_{1},\ldots,x_{d}]$ with $\deg f\lesssim_{d}A$ so that $f$ does not vanish identically on $V$ and each connected component of $\mathbb{R}^{d}\setminus Z(f)$ contains at most
$\lesssim_{d}\frac{|P|}{A^{i}\prod_{j=1}^{d-i}\delta_{j}(V)}$
points of $P$, where $\dim V\leq i\leq d$ is the unique integer with $\delta_{d-i}(V)\leq A<\delta_{d-i+1}(V)$.
As mentioned above, we can approximate any irreducible variety $V\subseteq\mathbb{C}^{d}$ by $Z(Q_{1},\ldots,Q_{d-\dim V})$ for some polynomials $Q_{1},\ldots,Q_{d-\dim V}$ with $\deg Q_{i}=\delta_{i}(V)$ (see [10, Lemma 5.3]). However, for technical reasons, we will need to make a slightly different choice for the polynomials $Q_{1},\ldots,Q_{d-\dim V}$. To be more specific, the other irreducible components of $Z(Q_{1},\ldots,Q_{d-\dim V})$ might have dimensions larger than expected, making it difficult to apply Theorem 2.2. Below we make a definition to identify those problematic components and then introduce another result of Walsh [10] that overcomes this issue.
Definition 3.4.
For an irreducible variety $V\subseteq\mathbb{C}^{d}$ of dimension $r$ and polynomials $Q_{1},\ldots,Q_{d-r}$ such that $V$ is an irreducible component of $Z(Q_{1},\ldots,Q_{d-r})$, we define the envelope $\mathcal{E}_{V}(Q_{1},\ldots,Q_{d-r})$ by
$\mathcal{E}_{V}(Q_{1},\ldots,Q_{d-r})=\bigcup_{i=1}^{d-r}\mathcal{E}_{V}^{(i)}(Q_{1},\ldots,Q_{d-r})$
where $\mathcal{E}_{V}^{(i)}(Q_{1},\ldots,Q_{d-r})$ is the union of the irreducible components of $Z(Q_{1},\ldots,Q_{i})$ of dimension at least $d-i+1$.
By Krull’s theorem, it is clear that $\dim_{x}^{\mathbb{C}}(Z(Q_{1},\ldots,Q_{d-r}))=r$ for all $x\in Z(Q_{1},\ldots,Q_{d-r})\setminus\mathcal{E}_{V}(Q_{1},\ldots,Q_{d-r})$. Therefore to make sure that we are in position to apply Theorem 2.2, we should manually excise $\mathcal{E}_{V}(Q_{1},\ldots,Q_{d-r})$ via a polynomial of small degree. This was done in [10, Section 6]. Expanding the definitions used in [10, Lemma 6.8] gives the following statement. (Actually it gives something strictly stronger, but we only state the parts that we will use.)
Proposition 3.5 ([10, Lemma 6.8]).
Given an irreducible variety $V\subseteq\mathbb{C}^{d}$ of dimension $j$, there exist polynomials $Q_{1},\ldots,Q_{d-j}$ and $F_{1},\ldots,F_{d-j-1}$ with the following properties:
  • $\deg Q_{k}\lesssim_{d}\delta_{k}(V)$ for each $1\leq k\leq d-j$;
  • $\deg F_{k}<\delta_{k}(V)$ for each $11\leq k\leq d-j-$ 1;
  • $V\subseteq Z(Q_{1},\ldots,Q_{d-j})$;
  • $F_{k}$ vanishes identically on each irreducible component of $\mathcal{E}_{V}(Q_{1},\ldots,Q_{d-j})$ of dimension $d-k$ for each $11\leq k\leq d-j-$ 1;
  • $F_{k}$ does not vanish identically on $V$ for each $11\leq k\leq d-j-$ 1.
We now prove Theorem 2.1 using Theorem 3.3 and Proposition 3.5.
Proof of Theorem 2.1.
We construct the desired data in steps numbered $d,d-1,\ldots,0$. At step $j$ we start with a set $\mathcal{U}_{j}$ whose elements are pairs $(V,P_{V})$ where $V\subseteq\mathbb{C}^{d}$ is an irreducible $j$-dimensional variety and $P_{V}\subseteq V\cap P$. During this step we will add some elements to $\mathcal{V}_{j},\mathcal{V}_{j+1},\ldots,\mathcal{V}_{d}$ and also produce a set $\mathcal{U}_{j-1}$.
To start, define $\mathcal{U}_{d}=\{(\mathbb{C}^{d},P)\}$. We will ensure that before step $j$, we have the partition
$P=\bigsqcup_{(V,P_{V})\in\mathcal{U}_{j}}P_{V}\sqcup\bigsqcup_{V\in\mathcal{V}}\bigsqcup_{C\in\mathcal{C}_{V}}P_{C}.$ (5)
Note that this is satisfied before step $d$.
Suppose we are given $\mathcal{U}_{j}$ for some $0\leq j\leq d$ and that 5 is satisfied. Step $j$ proceeds as follows. For each $(V,P_{V})\in\mathcal{U}_{j}$, define $j\leq i\leq d$ so that
$\delta_{1}(V)\leq\delta_{2}(V)\leq\cdots\leq\delta_{d-i}(V)\leq A<\delta_{d-i+1}(V)\leq\cdots\leq\delta_{d-j}(V).$
We apply Proposition 3.5 to $V$ to produce polynomials $Q_{1},\ldots,Q_{d-j}$ and $F_{1},\ldots,F_{d-j-1}$.
By real polynomial partitioning (Theorem 3.3) there exists a real polynomial $f_{\operatorname{part}}$ of degree $O_{d}(A)$ that does not vanish identically on $V$ such that each connected component of $\mathbb{R}^{d}\setminus Z(f_{\operatorname{part}})$ contains at most
$\lesssim_{d}\frac{\left\lvert P_{V}\right\rvert}{A^{i}\prod_{\ell=1}^{d-i}\delta_{\ell}(V)}$
points of $P_{V}$. Define $f_{V}=|F_{1}F_{2}\cdots F_{d-i-1}|^{2}f_{\operatorname{part}}$. We add $V=(Q^{(V)},f_{V},P_{V})$ to $\mathcal{V}_{i}$ where $Q^{(V)}=(Q_{1},\ldots,Q_{d-i})$. Note that this a $(d-i)$-tuple of polynomials; the last $i-j$ polynomials are not included in the data.
By our choice of $i$ and the above properties, we see that each of the polynomials $Q_{1},\ldots,Q_{d-i},f_{V}$ has degree $O_{d}(A)$. Clearly $f_{V}$ is a real polynomial. By definition, every irreducible component of $Z(Q_{1},\ldots,Q_{d-i})$ of dimension at least $i+1$ is contained in an irreducible component of $\mathcal{E}_{V}(Q_{1},\ldots,Q_{d-j})$ of dimension at least $i+1$, meaning that $f_{V}$ vanishes identically on it. Finally, $P_{V}\subseteq V\subseteq Z(Q_{1},\ldots,Q_{d-j})\subseteq Z(Q_{1},\ldots,Q_{d-i})$. This shows that $(Q^{(V)},f_{V},P_{V})$ has the desired properties.
We define $\mathcal{C}_{V}$ to be the set of connected components of $Z(Q^{(V)})(\mathbb{R})\setminus Z(f_{V})$ and for each $C\in\mathcal{C}_{V}$ we define $P_{C}=C\cap P_{V}$. Note that $\{P_{C}\}_{C\in\mathcal{C}_{V}}$ is a partition of $P_{V}\setminus Z(f_{V})$.
Let $W_{1},\ldots,W_{s}$ be the irreducible components of $V\cap Z(f_{V})$. Note that neither $f_{\operatorname{part}}$ nor any of $F_{1},\ldots,F_{d-i-1}$ vanish identically on $V$, meaning that $f_{V}$ does not vanish identically on $V$. Thus by Krull’s theorem, each $W_{k}$ has dimension $j-1$.
We define $P_{W_{k}}$ so that $P_{W_{k}}\subseteq W_{k}\cap P_{V}$ and so that they partition $P_{V}\cap Z(f_{V})$. This is clearly possible since $P_{V}\cap Z(f_{V})\subset V\cap Z(f_{V})=\bigcup_{k=1}^{s}W_{k}$. (There may be multiple ways to define these sets if some point lies in multiple of the $W_{k}$ but we make these choices arbitrarily.) We place $(W_{1},P_{W_{1}}),\ldots,(W_{s},P_{W_{s}})$ into $\mathcal{U}_{j-1}$. Note that since $\{P_{C}\}_{C\in\mathcal{C}_{V}}$ partitions $P_{V}\setminus Z(f_{V})$ and $\{P_{W_{k}}\}_{k\in[s]}$ partitions $P_{V}\cap Z(f_{V})$, we see that 5 is still satisfied. Doing this procedure for each $(V,\mathcal{P}_{V})\in\mathcal{V}_{j}$ completes step $j$.
We run this process for steps $j=d,d-1,\ldots,0$. At the end we have produced the desired data. For $V\in\mathcal{V}_{i}$, we chose $f_{\operatorname{part}}$ so that each connected component of $\mathbb{R}^{d}\setminus Z(f_{\operatorname{part}})$ contains at most
$\lesssim_{d}\frac{\left\lvert P_{V}\right\rvert}{A^{i}\prod_{j=1}^{d-i}\delta_{j}(V)}\lesssim_{d}\frac{\left\lvert P_{V}\right\rvert}{A^{i}\prod_{j=1}^{d-i}\deg Q_{j}^{(V)}}$
points of $P_{V}$. Since each connected components of $Z(Q^{(V)})(\mathbb{R})\setminus Z(f_{V})$ is contained in one of these connected components, we see that 4 is satisfied. By 5, we conclude that 3 is satisfied. Next note that for each $0\leq j\leq d$, the sets $\{P_{V}\}_{V\in\mathcal{A}_{j}}$ are disjoint, implying 2. Finally, since $\deg f_{V}\lesssim_{d}A$ for each $V$, by Bézout’s theorem, we see that for each $0\leq j\leq d$,
$\sum_{(V,P_{V})\in\mathcal{U}_{j}}\deg V\lesssim_{d}A^{d-j}.$
By Proposition 3.2, we have
$\deg V\gtrsim_{d}\delta_{1}(V)\cdots\delta_{d-j}(V)\gtrsim_{d}\delta_{1}(V)\cdots\delta_{d-i}(V)A^{j-i}\gtrsim_{d}\deg Q_{1}^{(V)}\cdots\deg Q_{d-i}^{(V)}A^{j-i}$
for every $(V,P_{V})\in\mathcal{U}_{j}$ with $i$ satisfying $\delta_{d-i}(V)\leq A<\delta_{d-i+1}(V)$. Combining the last two inequalities implies 1, proving the desired result. ∎
4. Turán results
In this section we prove Corollary 1.5. We deduce this from the following stronger result.
Proposition 4.1.
Let $H=(P_{1}\sqcup\cdots\sqcup P_{k},E)$ be a semialgebraic hypergraph in $\mathbb{R}^{d}$ with total degree $D$ and at least $\varepsilon|P_{1}|\cdots|P_{k}|$ edges. Fix $1\leq\ell\leq k$. There exist sets $S_{1}\subseteq P_{1},\ldots,S_{\ell}\subseteq P_{\ell}$ and $T\subseteq P_{\ell+1}\times\cdots\times P_{k}$ such that $S_{1}\times\cdots\times S_{\ell}\times T\subseteq E$ and such that for $1\leq i\leq\ell$,
$|S_{i}|=\Omega_{d,k}\left((\varepsilon/D)^{d}|P_{i}|\right)$
and
$|T|\geq\frac{\varepsilon}{3}|P_{\ell+1}|\cdots|P_{k}|.$
The $\ell=k-1$ of this result immediately implies Corollary 1.5. In Section 5 we will use the $\ell=1$ case in the derivation of Theorem 1.6.
Proof.
First we handle the $d\geq 2$ case. Fix $d,k,\ell$ and let $C=C(d,k)$ be a constant chosen sufficiently large in terms of the constant in Theorem 1.3. We will prove by induction on $|P_{1}|+\cdots+|P_{\ell}|$ that we can find the desired sets $S_{1},\ldots,S_{\ell},T$ satisfying
$|S_{i}|\geq\frac{1}{C}\left(\frac{\varepsilon}{D}\right)^{d}|P_{i}|$
for $1\leq i\leq\ell$ and
$|T|\geq\frac{\varepsilon}{3}|P_{\ell+1}|\cdots|P_{k}|.$
The base case $\left\lvert P_{1}\right\rvert+\cdots+\left\lvert P_{\ell}\right\rvert=0$ is trivial. For the inductive step, we apply Theorem 1.3 to $H$ with parameter $\varepsilon/3$ to produce partitions $\Pi_{1},\ldots,\Pi_{k}$. For each $1\leq i\leq\ell$, let $B_{i}$ be the union of parts $\pi\in\Pi_{i}$ with $\left\lvert\pi\right\rvert\leq\frac{1}{9k^{2}}\frac{|P_{i}|}{|\Pi_{i}|}$, i.e., $B_{i}$ is the union of the parts that are much smaller than average.
We split into two cases. If at most $(\varepsilon/3)\left\lvert P_{1}\right\rvert\cdots\left\lvert P_{k}\right\rvert$ edges involve a vertex in $B_{1}\cup\cdots\cup B_{\ell}$, define $H^{\prime}=(P_{1}\sqcup\cdots\sqcup P_{k},E^{\prime})$ to be the subhypergraph of $H$ formed by removing those edges along with the edges contained in an inhomogeneous $k$-tuple of parts. Then $|E^{\prime}|\geq(\varepsilon/3)\left\lvert P_{1}\right\rvert\cdots\left\lvert P_{k}\right\rvert$. By the pigeonhole principle, there exists some $\ell$-tuple $(x_{1},\ldots,x_{\ell})\in P_{1}\times\cdots\times P_{\ell}$ such that
$T=\{(x_{\ell+1},\ldots,x_{k})\in P_{\ell+1}\times\cdots\times P_{k}:(x_{1},\ldots,x_{k})\in E^{\prime}\}$
has size at least $(\varepsilon/3)|P_{\ell+1}|\cdots|P_{k}|$. For $1\leq i\leq\ell$, define $S_{i}\in\Pi_{i}$ to be the part of the partition containing $x_{i}$. We see that $S_{1}\times\cdots\times S_{\ell}\times T\subseteq E^{\prime}\subseteq E$ and for $1\leq i\leq\ell$,
$|S_{i}|\geq\frac{1}{9k^{2}}\frac{|P_{i}|}{|\Pi_{i}|}\gtrsim_{d,k}\left(\frac{\varepsilon}{D}\right)^{d}|P_{i}|.$
Choosing $C$ sufficiently large, this completes the induction in this case.
In the remaining case, there must be some $1\leq i\leq\ell$ such that the number of edges involving $B_{i}$ is at least $(\varepsilon/3k)\left\lvert P_{1}\right\rvert\cdots\left\lvert P_{k}\right\rvert$. Define $H^{\prime}=(P_{1}\sqcup\cdots\sqcup B_{i}\sqcup\cdots\sqcup P_{k},E^{\prime})$ to be the induced subhypergraph of $H$ on vertex set $P_{1}\sqcup\cdots\sqcup B_{i}\sqcup\cdots\sqcup P_{k}$. Note that
$|B_{i}|\leq\frac{1}{9k^{2}}\frac{|P_{i}|}{|\Pi_{i}|}|\Pi_{i}|=\frac{|P_{i}|}{9k^{2}}.$
Furthermore, the edge density of $H^{\prime}$ is
$\eta=\frac{|E^{\prime}|}{|P_{1}|\cdots|B_{i}|\cdots|P_{k}|}\geq\frac{\frac{\varepsilon}{3k}|P_{1}|\cdots|P_{k}|}{|P_{1}|\cdots|B_{i}|\cdots|P_{k}|}=\frac{\varepsilon}{3k}\cdot\frac{|P_{i}|}{|B_{i}|}\geq 3k\varepsilon>\varepsilon.$
Since $|P_{1}|+\cdots+|B_{i}|+\cdots+|P_{k}|<|P_{1}|+\cdots+|P_{k}|$, we can apply the inductive hypothesis to $H^{\prime}$. This produces $S_{1}\subseteq P_{1},\ldots,S_{i}\subseteq B_{i}\subset P_{i},\ldots,S_{\ell}\subseteq P_{\ell}$ and $T\subseteq P_{\ell+1}\times\cdots\times P_{k}$ such that $S_{1}\times\cdots\times S_{\ell}\times T\subseteq E^{\prime}\subseteq E$ and such that
$|S_{j}|\geq\frac{1}{C}\left(\frac{\eta}{D}\right)^{d}|P_{j}|\geq\frac{1}{C}\left(\frac{\varepsilon}{D}\right)^{d}|P_{j}|$
for each $1\leq j\leq\ell$ with $j\neq i$. Additionally,
$\begin{split}|S_{i}|\geq\frac{1}{C}\left(\frac{\eta}{D}\right)^{d}|B_{i}|&\geq\frac{1}{C}\left(\frac{\frac{\varepsilon}{3k}\cdot\frac{|P_{i}|}{|B_{i}|}}{D}\right)^{d}|B_{i}|=\frac{1}{C}\left(\frac{\varepsilon}{D}\right)^{d}\frac{\left(\frac{|P_{i}|}{|B_{i}|}\right)^{d-1}}{(3k)^{d}}|P_{i}|\\ &\geq\frac{1}{C}\left(\frac{\varepsilon}{D}\right)^{d}\frac{(9k^{2})^{d-1}}{(3k)^{d}}|P_{i}|\geq\frac{1}{C}\left(\frac{\varepsilon}{D}\right)^{d}|P_{i}|\end{split}$
where the last inequality holds since $d\geq 2$. Finally the inductive hypothesis also gives
$|T|\geq\frac{\eta}{3}|P_{\ell+1}|\cdots|P_{k}|\geq\frac{\varepsilon}{3}|P_{\ell+1}|\cdots|P_{k}|.$
Now we prove the result for $d=1$. Let $\Pi_{1},\ldots,\Pi_{k}$ be the equitable partitions formed by applying Corollary 1.4 with error parameter $\varepsilon/2$. Note that since $d=1$, we have $|\Pi_{i}|=O_{k}(D/\varepsilon)$. Let $H^{\prime}=(P_{1}\sqcup\cdots\sqcup P_{k},E^{\prime})$ be the subhypergraph of $H$ formed by removing the edges contained in an inhomegeneous $k$-tuple of parts. Clearly $|E^{\prime}|\geq(\varepsilon/2)\left\lvert P_{1}\right\rvert\cdots\left\lvert P_{k}\right\rvert$. By the pigeonhole principle, there exists some $\ell$-tuple $(x_{1},\ldots,x_{\ell})\in P_{1}\times\cdots\times P_{\ell}$ such that
$T=\{(x_{\ell+1},\ldots,x_{k})\in P_{\ell+1}\times\cdots\times P_{k}:(x_{1},\ldots,x_{k})\in E^{\prime}\}$
has size at least $(\varepsilon/2)|P_{\ell+1}|\cdots|P_{k}|$. For $1\leq i\leq\ell$, define $S_{i}\in\Pi_{i}$ to be the part of the partition containing $x_{i}$. We see that $S_{1}\times\cdots\times S_{\ell}\times T\subseteq E^{\prime}\subseteq E$ and that $S_{1},\ldots,S_{\ell},T$ are of the appropriate sizes. ∎
5. The Zarankiewicz problem
We prove the graph case ($k=2$) of Theorem 1.6 first. The proof is essentially a specialization of the general proof, but the calculations are significantly simpler so we think it is valuable to read this case first.
The graph case is proved in two steps. First we deduce a weak Zarankiewicz result from our Turán result (Corollary 1.5). Then we upgrade the weak Zarankiewicz result to the strong one via our multilevel polynomial partitioning scheme (Theorem 2.1). These are the same two steps used by Fox, Pach, Sheffer, Suk, and Zahl in the proof of their semialgebraic Zarankiewicz result. They prove a similar weak Zarankiewicz result [4, Theorem 2.1] using VC dimension techniques and then upgrade it using constant-degree polynomial partitioning. Note in contrast, both halves of our proof use only polynomial partitioning and we avoid losing $N^{o(1)}$ factors by using polynomial partitioning using polynomials of fairly high degree.
One should compare the following result to the Kővári–Sós–Turán theorem which gives the bound of $O_{u}(N^{2-1/u})$ edges. Here we improve this to $O_{u,d}(N^{2-1/d})$ which is typically much smaller.
Corollary 5.1.
Let $G=(P\sqcup Q,E)$ be a semialgebraic graph in $\mathbb{R}^{d}$ with total degree $D$. If $G$ is $K_{u,u}$-free, then the number of edges of $G$ is at most
$|E|=O_{d}\left(u^{1/d}D|P||Q|^{1-1/d}+u|Q|\right).$
Proof.
Let $|E|=\varepsilon|P||Q|$. By Corollary 1.5, $G$ contains $T\times S$ with
$|S|=\Omega_{d}\left((\varepsilon/D)^{d}|Q|\right)$
and
$|T|=\Omega(\varepsilon|P|).$
Since $G$ is $K_{u,u}$-free we conclude that one of these sets is smaller than $u$. Thus either
$(\varepsilon/D)^{d}|Q|=O_{d}(u)$
or
$\varepsilon|P|=O(u).$
The first inequality rearranges to
$|E|=\varepsilon|P||Q|=O_{d}\left(D\left(\frac{u}{|Q|}\right)^{1/d}|P||Q|\right)$
and the second to
$|E|=O(u|Q|).\qed$
Theorem 5.2.
Let $G=(P\sqcup Q,E)$ be a semialgebraic graph in $\mathbb{R}^{d}$ with total degree $D$. If $G$ is $K_{u,u}$-free, then the number of edges of $G$ is at most
$|E|=O_{d}\left(u^{\frac{2}{d+1}}D^{\frac{2d}{d+1}}(|P||Q|)^{\frac{d}{d+1}}+u|P|+u|Q|\right).$
Proof.
We first deal with the case that $u^{d-1}|Q|>D^{d-1}|P|^{d}$. In this case, Corollary 5.1 gives the bound
$|E|\lesssim_{d}u^{1/d}D|P||Q|^{1-1/d}+u|Q|.$
We will upper bound the first term by $u^{\frac{2}{d+1}}D(|P||Q|)^{\frac{d}{d+1}}$, which suffices as $1\leq\frac{2d}{d+1}$. To do so, simply note that
$u^{1/d}D|P||Q|^{1-1/d}=u^{1/d}D\left(\frac{\left\lvert P\right\rvert^{d}}{\left\lvert Q\right\rvert}\right)^{\frac{1}{d(d+1)}}(|P||Q|)^{\frac{d}{d+1}}<u^{\frac{2}{d+1}}D(|P||Q|)^{\frac{d}{d+1}}$
as $\left\lvert P\right\rvert^{d}/\left\lvert Q\right\rvert<u^{d-1}/D^{d-1}\leq u^{d-1}$ in this case.
Therefore we may now assume that $u^{d-1}\left\lvert Q\right\rvert\leq D^{d-1}\left\lvert P\right\rvert^{d}$. We apply the multilevel polynomial partitioning scheme (Theorem 2.1) to $P\subset\mathbb{R}^{d}$ with parameter
$A=u^{-1/(d+1)}D^{1/(d+1)}|P|^{d/(d^{2}-1)}|Q|^{-1/(d^{2}-1)}.$
Note that by assumption $A\geq 1$.
For each $q\in Q$, consider the semialgebraic neighborhood set
$N_{q}=\{x\in\mathbb{R}^{d}:\Phi(\operatorname{sgn}(g_{1}(x,q)),\ldots,\operatorname{sgn}(g_{t}(x,q)))=1\}.$
For $p\in P$ we have $(p,q)$ is an edge of $G$ if and only if $p\in N_{q}$.
For each cell $C\in\mathcal{C}$ we have a set $P_{C}$ produced by the theorem. Now for each cell, define a partition
$Q=Q_{C,\mathsf{none}}\sqcup Q_{C,\mathsf{some}}\sqcup Q_{C,\mathsf{all}}$
where $Q_{C,\mathsf{none}}=\{q\in Q:N_{q}\cap C=\emptyset\}$ and $Q_{C,\mathsf{all}}=\{q\in Q:N_{q}\cap C=C\}$ and the remaining vertices lie in $Q_{C,\mathsf{some}}$. Note in particular that $E[P_{C}\times Q_{C,\mathsf{none}}]$ is empty and $E[P_{C}\times Q_{C,\mathsf{all}}]$ is complete. Thus we can write
$|E|=\sum_{C\in\mathcal{C}}|E[P_{C}\times Q_{C,\mathsf{some}}]|+|E[P_{C}\times Q_{C,\mathsf{all}}]|.$
We bound the first sum using Corollary 5.1. For $V\in\mathcal{V}$, define $i_{V}(A)$ so that $V\in\mathcal{V}_{i_{V}(A)}$. Now
$\sum_{C\in\mathcal{C}}|E$ $[P_{C}\times Q_{C,\mathsf{some}}]|$
$=\sum_{V\in\mathcal{V}}\sum_{C\in\mathcal{C}_{V}}|E[P_{C}\times Q_{C,\mathsf{some}}]|$
$\lesssim_{d}\sum_{V\in\mathcal{V}}\sum_{C\in\mathcal{C}_{V}}u^{1/d}D|P_{C}||Q_{C,\mathsf{some}}|^{1-1/d}+u|Q_{C,\mathsf{some}}|$
$\lesssim_{d}\sum_{V\in\mathcal{V}}\left(u^{1/d}D\left(\frac{|P_{V}|}{A^{i_{V}(A)}\prod_{j=1}^{d-i_{V}(A)}\deg Q_{j}^{(V)}}\right)^{1-1/d}\sum_{C\in\mathcal{C}_{V}}|P_{C}|^{1/d}|Q_{C,\mathsf{some}}|^{1-1/d}\right.$
$\qquad\qquad\qquad\qquad\left.+u\sum_{C\in\mathcal{C}_{V}}|Q_{C,\mathsf{some}}|\right)$ [by 4]
$\leq\sum_{V\in\mathcal{V}}\left(u^{1/d}D|P_{V}|\left(\frac{\sum_{C\in\mathcal{C}_{V}}|Q_{C,\mathsf{some}}|}{A^{i_{V}(A)}\prod_{j=1}^{d-i_{V}(A)}\deg Q_{j}^{(V)}}\right)^{1-1/d}+u\sum_{C\in\mathcal{C}_{V}}|Q_{C,\mathsf{some}}|\right)$ [Hölder’s inequality]
For $q\in Q$ and $V\in\mathcal{V}$ we count the number of cells $C\in\mathcal{C}_{V}$ such that $q\in Q_{C,\mathsf{some}}$. If $N_{q}$ is neither entirely contained in nor entirely disjoint from $C$, we see that one of $\operatorname{sgn}(g_{1}(x,q)),\ldots,\operatorname{sgn}(g_{t}(x,q))$ is not constant on $C$. Thus by Lemma 2.4 (applied with $k=1$ and $g(x)=g_{r}(x,q)$), we see that $q\in Q_{C,\mathsf{some}}$ for at most $\lesssim_{d}DA^{i_{V}(A)-1}\prod_{j=1}^{d-i_{V}(A)}\deg Q_{j}^{(V)}$ choices of $C\in\mathcal{C}_{V}$. Summing over $q\in Q$ we conclude that
$\sum_{C\in\mathcal{C}_{V}}|Q_{C,\mathsf{some}}|\lesssim_{d}|Q|DA^{i_{V}(A)-1}\prod_{j=1}^{d-i_{V}(A)}\deg Q_{j}^{(V)}.$
Combining this with the above inequalities gives
$\sum_{C\in\mathcal{C}}|E[P_{C}\times Q_{C,\mathsf{some}}]|$ $\lesssim_{d}\sum_{V\in\mathcal{V}}\left(u^{1/d}D|P_{V}|\left(\frac{|Q|DA^{i_{V}(A)-1}\prod_{j=1}^{d-i_{V}(A)}\deg Q_{j}^{(V)}}{A^{i_{V}(A)}\prod_{j=1}^{d-i_{V}(A)}\deg Q_{j}^{(V)}}\right)^{1-1/d}\right.$
$\qquad\qquad\qquad\qquad\left.+u|Q|DA^{i_{V}(A)-1}\prod_{j=1}^{d-i_{V}(A)}\deg Q_{j}^{(V)}\right)$
$\lesssim_{d}\sum_{V\in\mathcal{V}}\left(u^{1/d}D^{2-1/d}|P_{V}||Q|^{1-1/d}A^{1/d-1}+u|Q|DA^{i_{V}(A)-1}\prod_{j=1}^{d-i_{V}(A)}\deg Q_{j}^{(V)}\right)$
$\lesssim_{d}u^{1/d}D^{2-1/d}|P||Q|^{1-1/d}A^{1/d-1}+u|Q|DA^{d-1}.$ [by 2 and 1]
By our choice of $A$, this is $O_{d}(u^{2/(d+1)}D^{2d/(d+1)}|P|^{d/(d+1)}|Q|^{d/(d+1)})$.
Now we bound the second sum. Note that since $G$ is $K_{u,u}$-free, we have either $|P_{C}|<u$ or $|Q_{C,\mathsf{all}}|<u$ for each $C\in\mathcal{C}$. We then have the easy bound
$\sum_{C\in\mathcal{C}:|P_{C}|\geq u}|E[P_{C}\times Q_{C,\mathsf{all}}]|$ $<\sum_{C\in\mathcal{C}:|P_{C}|\geq u}u|P_{C}|\lesssim_{d}u|P|.$
To finish, we define
$P^{\prime}=\bigcup_{C\in\mathcal{C}:|P_{C}|<u}P_{C}$
so that
$\sum_{C\in\mathcal{C}:|P_{C}|<u}|E[P_{C}\times Q_{C,\mathsf{all}}]|\leq|E[P^{\prime}\times Q]|\lesssim_{d}u^{1/d}D|P^{\prime}|^{1-1/d}|Q|+u|P^{\prime}|$
by Corollary 5.1.
We have the bound $|P^{\prime}|<u|\mathcal{C}|\lesssim_{d}uA^{d}$ by Lemma 2.3 together with 1. Using this bound and the trivial bound $|P^{\prime}|\leq|P|$, we see that the second sum is bounded by
$O_{d}\left(u^{1/d}D(uA^{d})^{1-1/d}|Q|+u|P|\right).$
As the first term simplifies to $u\left\lvert Q\right\rvert DA^{d-1}$, which has already appeared, the proof is complete. ∎
Now we give the general proof. The proof is by induction on the uniformity $k$. Given the result for $(k-1)$-uniform hypergraphs, we bootstrap it to a weak Zarankiewicz bound for $k$-uniform hypergraphs. This upgrading procedure uses our Turán result, specifically Proposition 4.1 for $\ell=1$. Then we upgrade the weak bound to the full theorem via the multilevel polynomial partitioning scheme.
Theorem 5.3.
Let $H=(P_{1}\sqcup\cdots\sqcup P_{k},E)$ be a semialgebraic hypergraph in $\mathbb{R}^{d}$ with total degree $D$. If $H$ is $K^{(k)}_{u,\ldots,u}$-free, then the number of edges of $H$ is at most
$|E|\lesssim_{d,k}\left\lvert P_{1}\right\rvert\cdots\left\lvert P_{k}\right\rvert\sum_{s=1}^{k}u^{\frac{s}{(s-1)d+1}}D^{\frac{(s-1)(s+2)d}{2\left((s-1)d+1\right)}}\sum_{I\in\binom{[k]}{s}}\left(\prod_{i\in I}\left\lvert P_{i}\right\rvert\right)^{-\frac{1}{(s-1)d+1}}.$
Proof.
We will prove the statement by induction on $k$. The statement when $k=1$ is trivial, and when $k=2$ the statement becomes Theorem 5.2. Now we assume that $k\geq 3$, and we assume that the statement holds for $k-1$.
As in the $k=2$ case, we will first need a weaker Zarankiewicz result. Suppose that there are $\varepsilon\left\lvert P_{1}\right\rvert\cdots\left\lvert P_{k}\right\rvert$ edges in $E$. By Proposition 4.1, there exists a set $S_{k}\subseteq P_{k}$ with size at least $\Omega_{d,k}\left((\varepsilon/D)^{d}\left\lvert P_{k}\right\rvert\right)$ and a set $T\subseteq P_{1}\times\cdots\times P_{k-1}$ with size at least $(\varepsilon/3)\left\lvert P_{1}\right\rvert\cdots\left\lvert P_{k-1}\right\rvert$ such that $S_{k}\times T\subseteq E$. Now if $\left\lvert S_{k}\right\rvert\geq u$ and there is a $K_{u,\ldots,u}^{(k-1)}$ in $T$, which we denote by $S_{1}\times\cdots\times S_{k-1}$, then $S_{1}\times\cdots\times S_{k}$ is a $K_{u,\ldots,u}^{(k)}$ in $H$. Therefore, we either have $\left\lvert S_{k}\right\rvert<u$ or, by the inductive hypothesis,
$|T|\lesssim_{d,k}\left\lvert P_{1}\right\rvert\cdots\left\lvert P_{k-1}\right\rvert\sum_{s=1}^{k-1}u^{\frac{s}{(s-1)d+1}}D^{\frac{(s-1)(s+2)d}{2\left((s-1)d+1\right)}}\sum_{I\in\binom{[k-1]}{s}}\left(\prod_{i\in I}\left\lvert P_{i}\right\rvert\right)^{-\frac{1}{(s-1)d+1}}.$
As a consequence, either $\varepsilon\lesssim_{d,k}u^{1/d}D\left\lvert P_{k}\right\rvert^{-1/d}$ or
$\varepsilon\lesssim_{d,k}\sum_{s=1}^{k-1}u^{\frac{s}{(s-1)d+1}}D^{\frac{(s-1)(s+2)d}{2\left((s-1)d+1\right)}}\sum_{I\in\binom{[k-1]}{s}}\left(\prod_{i\in I}\left\lvert P_{i}\right\rvert\right)^{-\frac{1}{(s-1)d+1}}.$
We can thus deduce that
$\frac{\left\lvert E\right\rvert}{\left\lvert P_{1}\right\rvert\cdots\left\lvert P_{k}\right\rvert}=\varepsilon\lesssim_{d,k}u^{1/d}D\left\lvert P_{k}\right\rvert^{-1/d}+\sum_{s=1}^{k-1}u^{\frac{s}{(s-1)d+1}}D^{\frac{(s-1)(s+2)d}{2\left((s-1)d+1\right)}}\sum_{I\in\binom{[k-1]}{s}}\left(\prod_{i\in I}\left\lvert P_{i}\right\rvert\right)^{-\frac{1}{(s-1)d+1}}.$
A quick note is that only the first term is potentially larger compared to the desired bound. This fact will be used later on.
After establishing a weak Zarankiewicz result using the inductive hypothesis, we now establish the inductive step. Apply Theorem 2.1 to $P_{1},\ldots,P_{k-1}$ with parameter $A\geq 1$, which produces data $\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(k-1)}$. For each $q\in P_{k}$ define the semialgebraic neighborhood $N_{q}\subseteq(\mathbb{R}^{d})^{k-1}$ to be
$N_{q}=\{x\in(\mathbb{R}^{d})^{k-1}:\Phi(\operatorname{sgn}(g_{1}(x,q)),\ldots,\operatorname{sgn}(g_{t}(x,q)))=1\}.$
For $V_{1}\in\mathcal{V}^{(1)},\ldots,V_{k-1}\in\mathcal{V}^{(k-1)}$ and each tuple of cells $C\in\mathcal{C}_{V_{1}}\times\cdots\times\mathcal{C}_{V_{k-1}}$, define the partition
$P_{k}=P_{C,\mathsf{none}}\sqcup P_{C,\mathsf{some}}\sqcup P_{C,\mathsf{all}}$
as before: $P_{C,\mathsf{none}}$ consists of points $q\in P_{k}$ with $N_{q}\cap C=\emptyset$, $P_{C,\mathsf{all}}$ consists of points $q\in P_{k}$ with $N_{q}\cap C=C$, and $P_{C,\mathsf{some}}$ consists of the remaining points in $P_{k}$. We get
$|E|=\sum_{C}|E[(P_{1})_{C_{1}}\times\cdots\times(P_{k-1})_{C_{k-1}}\times P_{C,\mathsf{some}}]|+|E[(P_{1})_{C_{1}}\times\cdots\times(P_{k-1})_{C_{k-1}}\times P_{C,\mathsf{all}}]|.$
We first deal with the first term. By the weak Zarankiewicz result, we may upper bound it by
$\sum_{C}u^{1/d}D\left\lvert(P_{1})_{C_{1}}\right\rvert\cdots\left\lvert(P_{k-1})_{C_{k-1}}\right\rvert\left\lvert P_{C,\mathsf{some}}\right\rvert^{1-1/d}$ (6)
$+$ $\sum_{C}\left\lvert(P_{1})_{C_{1}}\right\rvert\cdots\left\lvert(P_{k-1})_{C_{k-1}}\right\rvert\left\lvert P_{C,\mathsf{some}}\right\rvert\sum_{s=1}^{k-1}u^{\frac{s}{(s-1)d+1}}D^{\frac{(s-1)(s+2)d}{2\left((s-1)d+1\right)}}\sum_{I\in\binom{[k-1]}{s}}\left(\prod_{i\in I}\left\lvert(P_{i})_{C_{i}}\right\rvert\right)^{-\frac{1}{(s-1)d+1}}.$ (7)
We first fix a choice $(V_{1},\ldots,V_{k-1})$ and sum over $C\in\mathcal{C}_{V_{1}}\times\cdots\times\mathcal{C}_{V_{k-1}}$. Suppose that $V_{\ell}\in\mathcal{V}^{(\ell)}_{i_{\ell}}$ for each $\ell\in[k-1]$. By 4, we know that
$\left\lvert(P_{\ell})_{C_{\ell}}\right\rvert\lesssim_{d}\frac{\left\lvert(P_{\ell})_{V_{\ell}}\right\rvert}{A^{i_{\ell}}\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{i})}}.$
Also, as in the proof of Theorem 5.2, we still get from Lemma 2.4 that
$\sum_{C\in\mathcal{C}_{V_{1}}\times\cdots\times\mathcal{C}_{V_{k-1}}}|P_{C,\mathsf{some}}|\lesssim_{d,k}\left\lvert P_{k}\right\rvert DA^{-1}\prod_{\ell=1}^{k-1}\prod_{j=1}^{d-i_{\ell}(A)}A^{i_{\ell}}\deg Q_{j}^{(V_{\ell})}.$
Therefore summing over $C\in\mathcal{C}_{V_{1}}\times\cdots\times\mathcal{C}_{V_{k-1}}$, we may upper bound the term in 6 by
$\lesssim_{d}u^{1/d}D\left(\prod_{\ell=1}^{k-1}\frac{\left\lvert(P_{\ell})_{V_{\ell}}\right\rvert}{A^{i_{\ell}}\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{i})}}\right)^{1-1/d}\sum_{C\in\mathcal{C}_{V_{1}}\times\cdots\mathcal{C}_{V_{k-1}}}\left(\prod_{\ell=1}^{k-1}\left\lvert(P_{\ell})_{C_{\ell}}\right\rvert\right)^{1/d}\left\lvert P_{C,\mathsf{some}}\right\rvert^{1-1/d}.$
The sum may be upper bounded, using Hölder’s, by
$\left(\prod_{\ell=1}^{k-1}\left\lvert(P_{\ell})_{V_{\ell}}\right\rvert\right)^{1/d}\left(\left\lvert P_{k}\right\rvert DA^{-1}\prod_{\ell=1}^{k-1}\prod_{j=1}^{d-i_{\ell}}A^{i_{\ell}}\deg Q_{j}^{(V_{\ell})}\right)^{1-1/d}.$
Thus, when summing over $C\in\mathcal{C}_{V_{1}}\times\cdots\times\mathcal{C}_{V_{k-1}}$, the term in 6 can be upper bounded by
$O_{d,k}\left(u^{1/d}D^{2-1/d}A^{-1+1/d}\left\lvert(P_{1})_{V_{1}}\right\rvert\cdots\left\lvert(P_{k-1})_{V_{k-1}}\right\rvert\left\lvert P_{k}\right\rvert^{1-1/d}\right).$
Now sum this over $V_{1}\in\mathcal{V}^{(1)},\ldots,V_{k-1}\in\mathcal{V}^{(k-1)}$, and then we can bound 6 by
$O_{d,k}\left(u^{1/d}D^{2-1/d}A^{-1+1/d}\left\lvert P_{1}\right\rvert\cdots\left\lvert P_{k-1}\right\rvert\left\lvert P_{k}\right\rvert^{1-1/d}\right)$
using 2. Denote by $F(A)$ the quantity $u^{1/d}D^{2-1/d}A^{-1+1/d}\left\lvert P_{k}\right\rvert^{-1/d}$ accordingly.
Having bounded 6, we now turn to 7. Fix some $s\in[k-1]$ and some $I\in\binom{[k-1]}{s}$. We also fix $V_{1},\ldots,V_{k-1}$. Then for any $C\in\mathcal{C}_{V_{1}}\times\cdots\times\mathcal{C}_{V_{k-1}}$,
$u^{\frac{s}{(s-1)d+1}}D^{\frac{(s-1)(s+2)d}{2\left((s-1)d+1\right)}}\left\lvert P_{C,\mathsf{some}}\right\rvert\left(\prod_{\ell\in I}\left\lvert(P_{i})_{C_{i}}\right\rvert\right)^{1-\frac{1}{(s-1)d+1}}\prod_{\ell\in[k-1]\backslash I}\left\lvert(P_{i})_{C_{i}}\right\rvert$
$\lesssim_{d,k}$ $u^{\frac{s}{(s-1)d+1}}D^{\frac{(s-1)(s+2)d}{2\left((s-1)d+1\right)}}\left\lvert P_{C,\mathsf{some}}\right\rvert\left(\prod_{\ell\in I}\frac{\left\lvert(P_{\ell})_{V_{\ell}}\right\rvert}{A^{i_{\ell}}\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{i})}}\right)^{1-\frac{1}{(s-1)d+1}}\prod_{\ell\in[k-1]\backslash I}\frac{\left\lvert(P_{\ell})_{V_{\ell}}\right\rvert}{A^{i_{\ell}}\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{i})}}.$
Since we know that
$\sum_{C\in\mathcal{C}_{V_{1}}\times\cdots\times\mathcal{C}_{V_{k-1}}}|P_{C,\mathsf{some}}|\lesssim_{d,k}\left\lvert P_{k}\right\rvert DA^{-1}\prod_{\ell=1}^{k-1}\prod_{j=1}^{d-i_{\ell}(A)}A^{i_{\ell}}\deg Q_{j}^{(V_{\ell})},$
when we sum the term corresponding to $n$ and $I$ in 7 over $C\in\mathcal{C}_{V_{1}}\times\cdots\times\mathcal{C}_{V_{k-1}}$, we get
$\lesssim_{d,k}\left\lvert(P_{1})_{V_{1}}\right\rvert\cdots\left\lvert(P_{k})_{V_{k}}\right\rvert\left\lvert P_{k}\right\rvert D^{\frac{(s-1)(s+4)d+2}{2((s-1)d+1)}}A^{-1}\prod_{\ell\in I}\left(\frac{A^{i_{\ell}}\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{i})}}{\left\lvert(P_{\ell})_{V_{\ell}}\right\rvert}\right)^{\frac{1}{(s-1)d+1}}.$
For each $\ell\in I$, we know that
$\sum_{V_{\ell}\in\mathcal{V}^{(\ell)}}\left\lvert(P_{\ell})_{V_{\ell}}\right\rvert\left(\frac{A^{i_{\ell}}\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{i})}}{\left\lvert(P_{\ell})_{V_{\ell}}\right\rvert}\right)^{\frac{1}{(s-1)d+1}}$
$=$ $\sum_{V_{\ell}\in\mathcal{V}^{(\ell)}}\left\lvert(P_{\ell})_{V_{\ell}}\right\rvert^{1-\frac{1}{(s-1)d+1}}\left(A^{i_{\ell}}\prod_{j=1}^{d-i_{\ell}}\deg Q_{j}^{(V_{i})}\right)^{\frac{1}{(s-1)d+1}}$
$\lesssim_{d}$ $\left\lvert P_{\ell}\right\rvert^{1-\frac{1}{(s-1)d+1}}\left(A^{d}\right)^{\frac{1}{(s-1)d+1}}=\left\lvert P_{\ell}\right\rvert\left(\frac{A^{d}}{\left\lvert P_{\ell}\right\rvert}\right)^{\frac{1}{(s-1)d+1}}$
by Hölder’s inequality, 2 and 1. Therefore 7 can be upper bounded by
$O_{d,k}\left(\left\lvert P_{1}\right\rvert\cdots\left\lvert P_{k}\right\rvert\sum_{s=1}^{k-1}u^{\frac{s}{(s-1)d+1}}D^{\frac{(s-1)(s+4)d+2}{2((s-1)d+1)}}A^{\frac{d-1}{(s-1)d+1}}\sum_{I\in\binom{[k-1]}{s}}\prod_{\ell\in I}\left\lvert P_{\ell}\right\rvert^{-\frac{1}{(s-1)d+1}}\right).$
For each $I\subseteq[k-1]$ that is not empty, set $s=\left\lvert I\right\rvert$ and denote by $G_{I}(A)$ the quantity
$u^{\frac{s}{(s-1)d+1}}D^{\frac{(s-1)(s+4)d+2}{2((s-1)d+1)}}A^{\frac{d-1}{(s-1)d+1}}\prod_{\ell\in I}\left\lvert P_{\ell}\right\rvert^{-\frac{1}{(s-1)d+1}}.$
What we have proven is that for each $A\geq 1$,
$\sum_{C}|E[(P_{1})_{C_{1}}\times\cdots\times(P_{k-1})_{C_{k-1}}\times P_{C,\mathsf{some}}]|\lesssim_{d,k}F(A)+\sum_{\emptyset\neq I\subseteq[k-1]}G_{I}(A).$
It remains to bound $\sum_{C}|E[(P_{1})_{C_{1}}\times\cdots\times(P_{k-1})_{C_{k-1}}\times P_{C,\mathsf{all}}]|$. Note that for all $C$, either there exists $\ell\in[k-1]$ with $\left\lvert(P_{\ell})_{C_{\ell}}\right\rvert<u$ or $\left\lvert P_{C,\mathsf{all}}\right\rvert<u$. For the $C$’s with $\left\lvert P_{C,\mathsf{all}}\right\rvert<u$, it is clear that their contribution is at most $u\left\lvert P_{1}\right\rvert\cdots\left\lvert P_{k-1}\right\rvert$, which corresponds to a term in the claimed inequality (when $I=\{k\}$). For the remaining $C$’s, we may partition them into $k-1$ sets $\mathcal{B}_{1},\ldots,\mathcal{B}_{k-1}$ so that if $C\in\mathcal{B}_{\ell}$, then $\left\lvert(P_{\ell})_{C_{\ell}}\right\rvert<u$. Let $P^{\prime}_{\ell}$ be the union of $(P_{\ell})_{C_{\ell}}$’s with $\left\lvert(P_{\ell})_{C_{\ell}}\right\rvert<u$. As there are at most $O_{d}(A^{d})$ cells, we see that $\left\lvert P^{\prime}_{\ell}\right\rvert\lesssim_{d}uA^{d}$ for every $\ell\in[k-1]$. Moreover, the contributions of the $C$’s in $\mathcal{B}_{\ell}$ can be upper bounded by the number of edges involving $P_{\ell}^{\prime}$. By the weak Zarankiewicz we established earlier with $P_{\ell}^{\prime}$ being the last part, we get that
$\sum_{C\in\mathcal{B}_{\ell}}|E[(P_{1})_{C_{1}}\times\cdots\times(P_{k-1})_{C_{k-1}}\times P_{C,\mathsf{all}}]|$
$\lesssim_{d,k}$ $u^{1/d}D(uA^{d})^{1-1/d}\prod_{i\neq\ell}\left\lvert P_{i}\right\rvert+\prod_{i\in[k]}\left\lvert P_{i}\right\rvert\cdot\sum_{s=1}^{k-1}u^{\frac{s}{(s-1)d+1}}D^{\frac{(s-1)(s+2)d}{2\left((s-1)d+1\right)}}\sum_{I\in\binom{[k]\backslash\{\ell\}}{s}}\left(\prod_{i\in I}\left\lvert P_{i}\right\rvert\right)^{-\frac{1}{(s-1)d+1}}.$
Note that the second summand consists only of terms that appear in the claimed inequality. Therefore we just need to take care of the first summand, which simplifies to $uDA^{d-1}\prod_{i\neq\ell}\left\lvert P_{i}\right\rvert$. This is exactly $|P_{1}|\cdots|P_{k}|G_{\{\ell\}}(A)$. Therefore, if we can choose a positive integer $A$ such that
$F(A)+\sum_{\emptyset\neq I\subseteq[k-1]}G_{I}(A)\lesssim_{k}\sum_{s=1}^{k}u^{\frac{s}{(s-1)d+1}}D^{\frac{(s-1)(s+2)d}{2\left((s-1)d+1\right)}}\sum_{I\in\binom{[k]}{s}}\left(\prod_{i\in I}\left\lvert P_{i}\right\rvert\right)^{-\frac{1}{(s-1)d+1}},$ (8)
then we can close the induction.
Observe that $F$ is decreasing in $A$ while $G_{I}$ is increasing in $A$ for all $\emptyset\neq I\subseteq[k-1]$. For every $\emptyset\neq I\subseteq[k-1]$, set $M_{I}(A)$ to be the $A$ such that $F(A)=G_{I}(A)$. Note that if $\left\lvert I\right\rvert=s$, then
$F(M_{I}(A))=F(M_{I}(A))^{\frac{d}{sd+1}}G_{I}(M_{I}(A))^{\frac{(s-1)d+1}{sd+1}}=u^{\frac{s+1}{sd+1}}D^{\frac{s(s+3)d}{2(sd+1)}}\left(\prod_{i\in I\cup\{k\}}\left\lvert P_{i}\right\rvert\right)^{-\frac{1}{sd+1}},$
which is indeed a term in 8 that corresponds to $I\cup\{k\}$. Denote $\min_{\emptyset\neq I\subseteq[k-1]}M_{I}(A)$ by $m$. If $m\geq 1$, then we may choose $A=m$. By this choice of $A$, we have $G_{I}(A)\leq F(A)$ for all $\emptyset\neq I\subseteq[k-1]$, and we also have that $F(A)=F(m)$ is bounded by the right hand side of 8. Therefore, it remains to deal with the case where $m<1$. In this case, we can find a non-empty $I\subseteq[k-1]$ with $M_{I}(A)<1$. This shows that $F(1)<F(M_{I}(A))=G_{I}(M_{I}(A))<G_{I}(1)$ and so
$u^{1/d}D\left\lvert P_{k}\right\rvert^{-1/d}\leq F(1)<F(1)^{\frac{d}{sd+1}}G_{I}(1)^{\frac{(s-1)d+1}{sd+1}}=u^{\frac{s+1}{sd+1}}D^{\frac{s(s+3)d}{2(sd+1)}}\left(\prod_{i\in I\cup\{k\}}\left\lvert P_{i}\right\rvert\right)^{-\frac{1}{sd+1}},$
showing that the weak Zarankiewicz bound already suffices. ∎
6. Lower bound constructions
We begin with introducing some examples of semialgebraic graphs. We will gradually build on the examples and eventually obtain an example that shows the optimality of Theorem 1.3.
Example 6.1.
For every $m,d\geq 1$, let $P_{1}=P_{2}=[m]^{d}$. Define the graph $G=(P_{1}\sqcup P_{2},E)$ where $x$ is adjacent to $y$ if $x_{i}=y_{i}$ for some $i\in[d]$. This graph is semialgebraic (actually algebraic) in $\mathbb{R}^{d}$ of total degree $d$.
For any $\varepsilon>0$ and $m$ sufficiently large in terms of $d,\varepsilon$, we will show that any homogeneous partition of the above graph with error $\varepsilon$ must have $\Omega_{d}(\varepsilon^{-d})$ parts. This shows that the $\varepsilon$-dependence in Theorem 1.3 is optimal.
For our next examples, for each $k\geq 1$ we define real numbers $a_{0},\ldots,a_{k-1}\in[0,1)$ as follows. Let $p$ be the smallest prime greater than $k$. Then define $a_{i}\in[0,1)$ so that $i^{2}/p\equiv a_{i}\pmod{1}$.
Example 6.2.
For $D,m\geq 1$, let $a_{0},\ldots,a_{D-1}\in[0,1)$ be the real numbers defined above. Set $P_{1}=[m]$ and $P_{2}=[2m]$. Define the graph $G=(P_{1}\sqcup P_{2},E)$ where $x$ is adjacent to $y$ if an even number of the inequalities
$y-x-\frac{im}{D}-\frac{a_{i}m}{3D}\geq 0\qquad\text{for }0\leq i<D$
hold. This graph is semialgebraic in $\mathbb{R}^{1}$ of total degree $D$. Note that the graph can be either written in terms of $D$ polynomials of degree 1 or in terms of a single polynomial of degree $D$.
The edges in $G$ correspond to pairs $(x,y)$ lying in the black region in the following diagram.
For any $\varepsilon>0$, any $D\geq 1$, and any $m$ sufficiently large in terms of $D,\varepsilon$, we will show that any homogeneous partition of the above graph with error $\varepsilon$ must have $\Omega(D/\varepsilon)$ parts, showing that both the $\varepsilon$- and the $D$-dependencies in Theorem 1.3 are optimal when $d=1$. This also generalizes to higher dimensions.
Example 6.3.
For $D,m,d\geq 1$, set $k=\left\lfloor D/d\right\rfloor$. Let $a_{0},\ldots,a_{k-1}\in[0,1)$ be the real numbers defined above. Set $P_{1}=[m]^{d}$ and $P_{2}=[2m]^{d}$. Define the graph $G=(P_{1}\sqcup P_{2},E)$ where $x$ is adjacent to $y$ if an even number of the inequalities
$y_{j}-x_{j}-\frac{im}{D}-\frac{a_{i}m}{3D}\geq 0\qquad\text{for }0\leq i<k\text{ and }j\in[d]$
hold. This graph is semialgebraic in $\mathbb{R}^{d}$ of total degree $kd\leq D$. This graph can either be written in terms of $kd$ polynomials of degree 1 or in terms of a single polynomial of degree $kd$.
For any $\varepsilon>0$, any $D,d\geq 1$, and any $m$ sufficiently large in terms of $D,\varepsilon,d$, we will show that any homogeneous partition of the above graph with error $\varepsilon$ must have $\Omega_{d}((D/\varepsilon)^{d})$ parts, showing that both the $\varepsilon$- and the $D$-dependencies in Theorem 1.3 are optimal.
As the first step, we show that Example 6.1 indeed shows that the $\varepsilon$-dependence in Theorem 1.3 is optimal.
Proposition 6.4.
For every $d\geq 1$ and every $\varepsilon>0$, for $m$ sufficiently large, the graph in Example 6.1 has the property that any homogeneous partition with error $\varepsilon$ must have at least $(2\varepsilon)^{-d}$ parts.
Proof.
Suppose $m\geq d\varepsilon^{-(d+1)}$. Let $G$ be the graph in Example 6.1. For a set $A\subseteq P_{1}$, write $N(A)$ for the neighborhood of $A$ in $G$, i.e., the set of vertices in $P_{2}$ that are adjacent to at least one vertex in $A$. Also write $\pi_{j}(A)\subseteq[m]$ for the projection of $A$ onto the $j$th coordinate axis, i.e., the set of $a\in[m]$ such that there exists $x\in A$ with $x_{j}=a$.
Claim 6.5.
For $A\subseteq P_{1}$,
$|N(A)|\geq m^{d}-(m-|A|^{1/d})^{d}\geq|A|^{1/d}m^{d-1}.$
Proof.
For $y\in P_{2}$, we have that $y\not\in N(A)$ if and only if $y_{j}\not\in\pi_{j}(A)$ for all $j$. Thus we see that
$|N(A)|=m^{d}-\prod_{j=1}^{d}(m-|\pi_{j}(A)|)\geq m^{d}-\left(m-\frac{\sum_{j=1}^{d}|\pi_{j}(A)|}{d}\right)^{d}\geq m^{d}-\left(m-\left(\prod_{j=1}^{d}|\pi_{j}(A)|\right)^{1/d}\right)^{d}.$
by two applications of the AM-GM inequality.
Since $A\subseteq\pi_{1}(A)\times\cdots\times\pi_{d}(A)$, we see that $\prod_{j=1}^{d}|\pi_{j}(A)|\geq|A|$, so $|N(A)|\geq m^{d}-(m-|A|^{1/d})^{d}$. To deduce the last inequality in the lemma statement, note that
$1-\frac{|A|^{1/d}}{m}\geq\left(1-\frac{|A|^{1/d}}{m}\right)^{d}$
since the left-hand side is at most 1 and $d\geq 1$. Multiplying by $m^{d}$ and rearranging gives the desired inequality. ∎
Suppose for contradiction that $P_{1}=A_{1}\sqcup\cdots\sqcup A_{K}$ and $P_{2}=B_{1}\sqcup\cdots\sqcup B_{K}$ are partitions with $K<(2\varepsilon)^{-d}$ parts that form a homogeneous partition of $G$ with error $\varepsilon$. Now the graph $G$ is regular of degree $m^{d}-(m-1)^{d}\leq dm^{d-1}$. Therefore, if $E[A_{i}\times B_{j}]$ is complete, then $|A_{i}|,|B_{j}|\leq dm^{d-1}$. Thus
$\sum_{i,j:E[A_{i}\times B_{j}]\text{ non-empty}}\frac{|A_{i}||B_{j}|}{m^{2d}}\leq\varepsilon+\frac{K^{2}d^{2}}{m^{2}}\leq 2\varepsilon,$
where the last inequality follows from the assumption that $m$ is sufficiently large. Furthermore, we can bound
$\sum_{i,j:E[A_{i}\times B_{j}]\text{ non-empty}}\frac{|A_{i}||B_{j}|}{m^{2d}}$ $\geq\sum_{i=1}^{K}\frac{|A_{i}||N(A_{i})|}{m^{2d}}$
$\geq\sum_{i=1}^{K}\frac{|A_{i}|^{1+\tfrac{1}{d}}m^{d-1}}{m^{2d}}$ [6.5]
$=\frac{K}{m^{d+1}}\left(\frac{1}{K}\sum_{i=1}^{K}|A_{i}|^{1+\tfrac{1}{d}}\right)$
$\geq\frac{K}{m^{d+1}}\left(\frac{1}{K}\sum_{i=1}^{K}|A_{i}|\right)^{1+\tfrac{1}{d}}$ [Jensen’s inequality]
$=\frac{1}{K^{1/d}}$
Combining this lower bound with the previous upper bound, we conclude that $K\geq(2\varepsilon)^{-d}$, as desired. ∎
We see that understanding the behavior of neighborhoods plays an important role in the proof. In the proof that both the $\varepsilon$- and the $D$-dependencies are optimal, we will show that neighborhoods of far-away points differ a lot, which will again be key in the proof.
Proposition 6.6.
For every $D\geq d\geq 1$ and every $\varepsilon\in(0,1/400)$, for $m$ sufficiently large the graph in Example 6.3 has the property that any homogeneous partition with error $\varepsilon$ must have at least $(D/(16\varepsilon d))^{d}$ parts.
Proof.
Suppose $m\geq 2400D\varepsilon^{-1}$. Let $G$ be the graph in Example 6.3 with parameters $D,m,d$. Recall that $k=\left\lfloor D/d\right\rfloor$. Let $H$ be the graph in Example 6.2 with parameters $k,m$. We first show that the numbers $a_{0},\ldots,a_{k-1}$ in the definitions of $G$ and $H$ are well-spaced in the following way.
Claim 6.7.
For all $1\leq\ell\leq k$, the number of pairs $(i,i^{\prime})\in\{0,1\ldots,k-\ell-1\}^{2}$ satisfying
$|a_{i+\ell}-a_{i}-a_{i^{\prime}+\ell}+a_{i^{\prime}}|<\frac{1}{16}$
is at most $k^{2}/4$.
Proof.
First note that $k<p\leq 2k$ since $p$ is the smallest prime greater than $k$. Now if $(i,i^{\prime})$ satisfies the above condition, then
$a_{i+\ell}-a_{i}-a_{i^{\prime}+\ell}+a_{i^{\prime}}\equiv\frac{(i+\ell)^{2}-i^{2}-(i^{\prime}+\ell)^{2}+i^{\prime 2}}{p}\equiv\frac{2\ell(i-i^{\prime})}{p}\pmod{1}$
is within $1/16$ of an integer. Since $\ell\not\equiv 0\pmod{p}$, this implies that there are at most $p/8$ possible values of $i-i^{\prime}$ and thus at most $kp/8\leq k^{2}/4$ possible values for $(i,i^{\prime})$. ∎
We next show that vertices that are far apart have very different neighborhoods in $H$ and also that these neighborhoods are not close to complementary.
Claim 6.8.
For $x,x^{\prime}\in[m]$,
$|N_{H}(x)\Delta N_{H}(x^{\prime})|,2m-|N_{H}(x)\Delta N_{H}(x^{\prime})|\geq\min\left\{\frac{k(|x-x^{\prime}|-1)}{2},\frac{m}{200}\right\}.$
Proof.
Define the two sets
$S=\left\{x+\frac{im}{k}+\frac{a_{i}m}{3k}:0\leq i<k\right\},$
and
$T=\left\{x^{\prime}+\frac{im}{k}+\frac{a_{i}m}{3k}:0\leq i<k\right\}.$
Since $a_{i}\in[0,1)$, the elements of $S$ are each separated by at least $2m/3k$. The same holds for the elements of $T$. We claim that there are at most $k/2$ pairs $(i,i^{\prime})\in\{0,\ldots,k-1\}^{2}$ such that the $i$-th element of $S$ and the $i^{\prime}$-th element of $T$ are at distance smaller than $\min\{|x-x^{\prime}|,m/(96k)\}$. For such a pair $(i,i^{\prime})$, we have
$\left\lvert\left(x+\frac{im}{k}+\frac{a_{i}m}{3k}\right)-\left(x^{\prime}+\frac{i^{\prime}m}{k}+\frac{a_{i^{\prime}}m}{3k}\right)\right\rvert<\min\left\{|x-x^{\prime}|,\frac{m}{96k}\right\}.$ (9)
Note that if $i=i^{\prime}$, the above inequality would imply that $|x-x^{\prime}|<|x-x^{\prime}|$ which does not hold. Thus we have $i\neq i^{\prime}$.
Now 9 rearranges to the containment
$\frac{k}{m}(x-x^{\prime})\in\left(i^{\prime}-i+\frac{a_{i^{\prime}}-a_{i}}{3}-\frac{1}{96},i^{\prime}-i+\frac{a_{i^{\prime}}-a_{i}}{3}+\frac{1}{96}\right)\subset\left(i^{\prime}-i-\frac{1}{2},i^{\prime}-i+\frac{1}{2}\right).$
Thus we see that $i^{\prime}-i=\ell\neq 0$ where $\ell$ is the closest integer to $(k/m)(x-x^{\prime})$. Thus the original inequality implies that
$\left\lvert x-x^{\prime}-\frac{\ell m}{k}-\frac{(a_{i+\ell}-a_{i})m}{3k}\right\rvert<\frac{m}{96k}.$
If $(i,i+\ell)$ and $(i^{\prime},i^{\prime}+\ell)$ both satisfy 9, we then have
$|a_{i+\ell}-a_{i}-a_{i^{\prime}+\ell}+a_{i^{\prime}}|\leq\frac{3k}{m}\left\lvert x-x^{\prime}-\frac{\ell m}{k}-\frac{(a_{i+\ell}-a_{i})m}{3k}\right\rvert-\frac{3k}{m}\left\lvert x-x^{\prime}-\frac{\ell m}{k}-\frac{(a_{i^{\prime}+\ell}-a_{i}^{\prime})m}{3k}\right\rvert<\frac{1}{16}.$
Therefore by 6.7, we see that there are at most $k/2$ pairs of elements of $S\sqcup T$ that are at distance smaller than $\min\{|x-x^{\prime}|,m/(96k)\}$.
The $2k$ elements of $S\sqcup T$ divide the interval $[0,2m]$ into $12k+$ 1 subintervals. The first quantity, $|N_{H}(x)\Delta N_{H}(x^{\prime})|$, is the number of points of $[2m]$ which lie in the even numbered subintervals. The second quantity, $2m-|N_{H}(x)\Delta N_{H}(x^{\prime})|$ is the number of points which lie in the odd numbered subintervals. Thus we are summing the number of integer points in at least $k$ subintervals. Of all the subintervals, we have shown that at most $k/2$ have length strictly smaller than $\min\{|x-x^{\prime}|,m/(96k)\}$ so we end up summing at least $k/2$ subintervals, each containing at least $\min\{|x-x^{\prime}|,m/(96k)\}-1$ integer points. Therefore we have shown that
$|N_{H}(x)\Delta N_{H}(x^{\prime})|\geq\frac{k}{2}\left(\min\left\{|x-x^{\prime}|,\frac{m}{96k}\right\}-1\right).$
This proves the desired inequality since $m\geq 2400k$. ∎
Now we claim that for $x,x^{\prime}\in[m]^{d}$ we have
$|N_{G}(x)\Delta N_{G}(x^{\prime})|\geq(2m)^{d-1}\cdot\min\left\{\frac{k(\|x-x^{\prime}\|_{\infty}-1)}{2},\frac{m}{200}\right\}.$ (10)
To see this, fix $j\in[d]$ such that $|x_{j}-x^{\prime}_{j}|=\|x-x^{\prime}\|_{\infty}$. For any $z\in[2m]^{[d]\setminus\{j\}}$, consider the line $L_{z}=\{y\in[2m]^{d}:y_{i}=z_{i}\text{ for all }i\in[d]\setminus\{j\}\}$. Clearly these $(2m)^{d-1}$ lines partition $P_{2}$. Furthermore, it follows from the definition of $G$ and $H$ that for each $z$ either $L_{z}\cap N_{G}(x)$ is $N_{H}(x_{j})$ or its complement. Now by 6.8 we see that $L_{z}\cap(N_{G}(x)\Delta N_{G}(x^{\prime}))$ is large for each $z$. Summing over $z$ gives 10.
To complete the proof, suppose for contradiction that $P_{1}=A_{1}\sqcup\cdots\sqcup A_{K}$ and $P_{2}=B_{1}\sqcup\cdots\sqcup B_{K}$ are partition with $K\leq(D/(16d\varepsilon))^{-d}$ parts that form a homogeneous partition of $G$ with error $\varepsilon$. Pick points $x_{i},x^{\prime}_{i}\in A^{\prime}_{i}$ that achieve $\max_{x_{i},x^{\prime}_{i}\in A_{i}}\|x_{i}-x^{\prime}_{i}\|_{\infty}$. Note that $|A_{i}|\leq(\|x^{\prime}_{i}-x_{i}\|_{\infty}+1)^{d}$ and also for any part $B_{j}$ with $y\in B_{j}$ such that $y\in N_{G}(x_{i})\Delta N_{G}(x_{i}^{\prime})$, we have that $E[A_{i}\times B_{j}]$ is not homogeneous. Thus by 10,
$\varepsilon$ $\geq\sum_{i,j:E[A_{i}\times B_{j}]\text{ not homogeneous}}\frac{|A_{i}||B_{j}|}{m^{d}(2m)^{d}}$
$\geq\sum_{i}\frac{|A_{i}||N_{G}(x_{i})\Delta N_{G}(x^{\prime}_{i})|}{m^{d}(2m)^{d}}$
$\geq\sum_{i}\min\left\{\frac{(2m)^{d-1}|A_{i}|k(\|x^{\prime}_{i}-x_{i}\|_{\infty}-1)}{2m^{d}(2m)^{d}},\frac{(2m)^{d-1}|A_{i}|m}{200m^{d}(2m)^{d}}\right\}$
$\geq\sum_{i}\min\left\{\frac{k|A_{i}|(|A_{i}|^{1/d}-2)}{4m^{d+1}},\frac{|A_{i}|}{400m^{d}}\right\}$
$=\min\left\{\frac{k}{4}\sum_{i}\frac{|A_{i}|^{\tfrac{d+1}{d}}}{m^{d+1}}-\frac{k}{2m}\sum_{i}\frac{|A_{i}|}{m^{d}},\frac{1}{400}\sum_{i}\frac{|A_{i}|}{m^{d}}\right\}$
$\geq\min\left\{\frac{k}{4K^{1/d}}-\frac{k}{2m},\frac{1}{400}\right\},$
where we apply Jensen’s inequality once again. Since $\varepsilon<1/400$ and $m$ is sufficiently large, we conclude that $\varepsilon\geq k/(8K^{1/d})$. Since $k\geq D/(2d)$, this rearranges to the desired result. ∎
Appendix A Deduction of polynomial method results
A.1. Proof of Theorem 2.2
For a variety $V\subseteq\mathbb{C}^{d}$ and a point $x\in V(\mathbb{R})$, we write $\dim_{x}^{\mathbb{R}}(V)$ for the local real dimension of $V(\mathbb{R})$ at $x$. Clearly $\dim_{x}^{\mathbb{R}}(V)\leq\dim_{x}^{\mathbb{C}}(V)$ for every $x\in V(\mathbb{R})$. We will need the following definitions that are made in [10, Section 7.1].
Definition A.1.
Let $Q=(Q_{1},Q_{2},\ldots,Q_{m})$ be a sequence of polynomials in some ambient space $\mathbb{C}^{d}$. For any $0\leq j\leq m$, define $Z_{j}(Q)$ to be $Z(Q_{1},\ldots,Q_{j})$, and for every $x\in\mathbb{R}^{d}$ let $\dim^{\mathbb{R}}_{Q,(j)}(x)$ be the $j$-tuple recording the local real dimensions $(\dim^{\mathbb{R}}_{x}Z_{1}(Q),\ldots,\dim^{\mathbb{R}}_{x}Z_{j}(Q))$. Lastly, for any $j$-tuple of non-negative integers $\tau$, define $Z_{\tau}(Q)$ to be the set
$\overline{\{x\in\mathbb{R}^{d}:\dim^{\mathbb{R}}_{Q,(j)}(x)=\tau\}}$
where the closure is in the Euclidean topology on $\mathbb{R}^{d}$.
Proof of Theorem 2.2.
We first reorder $Q_{1},\ldots,Q_{d-r}$ so that their degrees are non-decreasing. Now for every $x\in Z(Q)(\mathbb{R})$ with $\dim_{x}^{\mathbb{C}}Z(Q)=r$, by Krull’s theorem, we must have that $\dim_{x}^{\mathbb{C}}Z_{j}(Q)=d-j$ for every $j=1,\ldots,d-r$. As a consequence, we have $\dim^{\mathbb{R}}_{Q,(m)}(x)\leq(d-1,d-2,\ldots,r)$. Therefore we may now apply [10, Proposition 7.1] with $j=d-r$ and sum over all $d-r$-tuples $\tau$ with $\tau_{j}\leq d-j$ for all $j\in[d-r]$ to get the desired bound. ∎
A.2. Proof of Theorem 3.3
We deduce Theorem 3.3 from [10, Theorem 3.2]. The statement of that theorem involves the quantities $\Delta_{i}(V)$ and $i_{V}(A)$. We refer the reader to [10, Section 3] for the definition of $i_{V}(A)$ since our proof does not use any property of $i_{V}(A)$ other than that it is an integer between 0 and $d-\dim V$.
Let $V\subseteq\mathbb{C}^{d}$ be an irreducible variety. For every index $0\leq i\leq d-\dim V$, write
$\Delta_{i}(V)=\max\left(\frac{\deg V}{\delta_{i+1}(V)\cdots\delta_{d-\dim V}(V)},1\right).$
Note that in particular, we have $\Delta_{d-\dim V}(V)=\deg V$, and so the middle term in Proposition 3.2 can be rewritten as $\Delta_{d-\dim V}$. We can extend Proposition 3.2 as follows.
Proposition A.2.
For any irreducible variety $V\subseteq\mathbb{C}^{d}$ and any index $0\leq i\leq d-\dim V$,
$\prod_{j=1}^{i}\delta_{j}(V)\lesssim_{d}\Delta_{i}(V)\leq\prod_{j=1}^{i}\delta_{j}(V).$
Proof.
By Proposition 3.2,
$\Delta_{i}(V)\geq\frac{\deg V}{\prod_{j=i+1}^{d-\dim V}\delta_{j}(V)}\gtrsim_{d}\frac{\prod_{j=1}^{d-\dim V}\delta_{j}(V)}{\prod_{j=i+1}^{d-\dim V}\delta_{j}(V)}=\prod_{j=1}^{i}\delta_{j}(V).$
For the other direction, it clearly holds if $\Delta_{i}(V)=1$. Otherwise, by Proposition 3.2,
$\Delta_{i}(V)=\frac{\deg V}{\prod_{j=i+1}^{d-\dim V}\delta_{j}(V)}\leq\frac{\prod_{j=1}^{d-\dim V}\delta_{j}(V)}{\prod_{j=i+1}^{d-\dim V}\delta_{j}(V)}=\prod_{j=1}^{i}\delta_{j}(V).\qed$
Proof of Theorem 3.3.
By directly applying [10, Theorem 3.2], it remains to show that
$\frac{\left\lvert P\right\rvert}{A^{d-i_{V}(A)}\Delta_{i_{V}(A)}(V)}\lesssim_{d}\frac{\left\lvert P\right\rvert}{A^{i}\prod_{j=1}^{d-i}\delta_{j}(V)}.$
Since $\Delta_{i_{V}(A)}(V)\gtrsim_{d}\prod_{j=1}^{i_{V}(A)}\delta_{j}(V)$ by Proposition A.2, it suffices to show that
$A^{d-i_{V}(A)}\prod_{j=1}^{i_{V}(A)}\delta_{j}(V)\geq A^{i}\prod_{j=1}^{d-i}\delta_{j}(V).$
However, by the definition of $i$ and the fact that $\delta_{1}(V)\leq\cdots\leq\delta_{d-\dim V}(V)$, we know that
$A^{i}\prod_{j=1}^{d-i}\delta_{j}(V)=\min_{\dim V\leq\ell\leq d}\left\{A^{\ell}\prod_{j=1}^{d-\ell}\delta_{j}(V)\right\}.$
Therefore the inequality follows. ∎